### v1

# EquiContact: A Hierarchical SE(3) Vision-to-Force Equivariant Policy for Spatially Generalizable Contact-rich Tasks

Abstract
This paper presents a framework for learning vision-based robotic policies for contact-rich manipulation tasks that generalize spatially across task configurations. We focus on achieving robust spatial generalization of the policy for the peg-in-hole (PiH) task trained from a small number of demonstrations. We propose EquiContact, a hierarchical policy composed of a high-level vision planner (Diffusion Equivariant Descriptor Field, Diff-EDF) and a novel low-level compliant visuomotor policy (Geometric Compliant ACT, G-CompACT). G-CompACT operates using only localized observations (geometrically consistent error vectors (GCEV), force-torque readings, and wrist-mounted RGB images) and produces actions defined in the end-effector frame. Through these design choices, we show that the entire EquiContact pipeline is $SE(3)$-equivariant, from perception to force control. We also outline three key components for spatially generalizable contact-rich policies: compliance, localized policies, and induced equivariance. Real-world experiments on PiH tasks demonstrate a near-perfect success rate and robust generalization to unseen spatial configurations, validating the proposed framework and principles. The experimental videos can be found on the project website: https://sites.google.com/berkeley.edu/equicontact
I Introduction
Imitation learning has recently shown significant success in expanding the capabilities of machine learning in real-world robotics applications [1, 2]. In the early stages of robot learning, many methods formulated manipulation as sequences of keyframe-based pick-and-place actions [3, 4]. More works have started to produce a continuous set of actions directly from vision inputs [5, 6, 7]. Similar to the trend seen in large language models (LLMs), there is a growing belief that large-scale data can unlock generalizable, vision-based policies for robotics [8]. This has led to massive efforts to build large datasets [1] for training policies with general knowledge.
However, such policies often lack spatial generalizability and therefore require a large amount of data to learn robust policies. As described in [9], both action chunking transformers (ACT) [6] and diffusion policy (DP) [5] exhibit near-linear performance growth as the demonstration dataset size increases, suggesting that the trained policies do not inherently generalize well to new spatial configurations.
An alternative line of recent research focuses on leveraging symmetry—particularly equivariance—to enhance spatial generalizability, thereby improving sample efficiency during training [10, 11]. This approach requires fewer data but comes with its own challenges. Equivariant neural networks, being of a more specialized nature, are often not as well-developed and are more computationally intensive than their non-equivariant counterparts, making real-time and large-scale deployment more difficult. As a result, it becomes more attractive for users to use standard models trained with massive datasets in many instances.
[…]
In this paper, we propose EquiContact, a hierarchical SE(3) vision-to-force equivariant policy for spatially generalizable contact-rich tasks. Our proposed method consists of two main components. The first is a high-level planner consisting of a Diff-EDF model [14], which uses global point cloud data to provide a local reference frame for placing the peg relative to the hole. The second, low-level compliant visuomotor policy is a variant of ACT [6], which we refer to as Geometric Compliant ACT (G-CompACT). This policy handles detailed motion and contact interaction using force-torque feedback and RGB inputs from wrist-mounted cameras. A key design feature of G-CompACT is that it only relies on local information—specifically, the force-torque signal in the end-effector frame, a geometrically consistent error vector (GCEV) [11], and wrist camera inputs. The output of G-CompACT is the desired pose and admittance gains, which are then sent to the geometric admittance controller (GAC) module to execute compliant control. We emphasize that our contribution is on the structural framework rather than the specific choice of algorithms. For example, one could replace the Diff-EDF with other equivariant methods, such as ET-SEED [15], or replace the ACT with the DP or their variants, e.g., Diffusion Transformer (DiT) [16]. The main contributions of this paper are as follows:
  1. 1.
    We propose EquiContact, a hierarchical, provably $SE(3)$-equivariant policy from point clouds and RGB inputs to interaction forces for executing contact-rich tasks.
  2. 2.
    We demonstrate that EquiContact achieves near-perfect success rates and spatial generalizability in real robot experiments involving tight insertions.
  3. 3.
    We identify three key principles for spatially generalizable contact-rich manipulation: Compliance, Localized Policy (Left invariance), and Induced Equivariance. These enable $SE(3)$-equivariant behavior without requiring explicitly equivariant neural networks [10].
From these key principles, we propose a general framework to enhance the spatial generalization and interpretability of vision-based policies, namely, “anchoring localized policy on globally estimated reference frame.” We emphasize that our work provides complementary insights to recent trends of robot learning [1, 8, 16, 2] that aim to build generalist policies from a large-scale demo dataset. Our principles offer structural guidelines to improve spatial generalizability through $SE(3)$ equivariance.
II Related Works
Visuomotor Servoing Methods Recently, generative modeling has become mainstream in realizing visuomotor servoing policies. Particularly, there are two dominant methods for visuomotor servoing: Action Chunking with Transformers (ACT) [6] and Diffusion Policy (DP) [5]. ACT utilizes the conditional variational autoencoder (CVAE) method as a generative model, whereas DP employs denoising diffusion as its generative model. ACT and DP have been extended to other approaches, including compliance and force-reactive behaviors [17, 18, 19], as well as structural improvements [20, 16, 21]. Our work is most closely related to CompliantACT (CompACT) [17], which integrates compliant control for visuomotor policies. We have significantly improved CompACT by incorporating a provable $SE(3)$ equivariant structure.
Equivariant Methods Equivariant methods aim to leverage the inherent symmetries of robot tasks, typically modeled as group transformations such as $SE(3)$, to improve the sample efficiency of imitation learning. Early approaches, such as Neural Descriptor Fields (NDFs) [22] and Transporter Networks [3], incorporated equivariance to $SE(3)$ and $SO(2)$ transformations, utilizing scene point clouds and top-down views, respectively. Equivariant Descriptor Fields (EDFs) [23] reformulated the NDF approach into a fully end-to-end learning method with $SE(3)$ bi-equivariance property, which is further improved by Diffusion-EDF (Diff-EDF) [14] to improve training and inference performance. Other equivariant models that extract desired keyframes from point clouds include [24]. 3D Equivariant extensions of DP [25, 26, 15] and Flow-Matching [27, 28] policies have been explored in recent literature and combine $SE(3)$ equivariant observation encoders with appropriately modified diffusion/flow-matching processes. Compared to recent works of [29, 30], we generalize the table-top manipulation to contact-rich tasks while achieving full $SE(3)$ equivariance from vision to force/control level, bridging vision, control, and force interaction under a unified framework.
Figure 1: We propose an EquiContact, a hierarchical, provably $SE(3)$ vision-to-force equivariant policy for spatially generalizable contact-rich tasks. The proposed EquiContact consists of G-CompACT and Diffusion-EDF. The G-Compact plays a localized policy over the reference frame provided by the Diffusion-EDF, making our framework generalizable to unseen scenarios during evaluation.
III Problem Definition
In this paper, we aim to identify the key structural components of a policy capable of executing spatially generalizable, contact-rich manipulation tasks. As mentioned earlier, our proposed building blocks to achieve $SE(3)$ vision-to-force equivariant policy are: 1. Compliant Control, 2. Localized Policy, and 3. Induced Equivariance - which we will experimentally validate through specific data collection and evaluation setups, as detailed below.
III-A Problem Setup
Our primary focus is to provide a learning-based solution for a peg-in-hole task,a classic representative of contact-rich, force-based robotic manipulation and assembly tasks, using vision. Unlike prior work such as [11], which assumes that the desired hole pose is known and that the robot starts with a pre-grasped peg, we consider a more general scenario. The robot first needs to pick up the peg and then insert it into the hole, as shown in Fig. 2, relying solely on vision and proprioception. We assume only that the peg is upright in the scene. Following the reasoning that every manipulation task can be regarded as a sequence of pick-and-place, we formulated peg-in-hole insertion as a compliant pick-and-place task. However, as highlighted in the earlier discussion, naively using the pick-and-place approach is insufficient because the high-level vision planner often lacks the accuracy required for precise and contact-rich tasks, which is $1\mathrm{mm}$ in our case (see Fig. 2).
To accommodate this, we propose a low-level policy that:
  • provides real-time visual feedback to correct errors from the coarse high-level planner,
  • handles fine force-based interactions through compliance,
  • and achieves provable $SE(3)$-equivariance.
Assuming that a high-level vision planner that returns approximate desired end-effector pick-and-place poses is readily available, we focus on developing equivariant compliant placing, i.e., an insertion policy via imitation learning, and regard picking as a spatial case of placing.
To train this policy, we collect expert demonstrations of insertions on a fixed-platform setting with a known hole location. This enables the policy to perform nearly perfectly in the trained scenario. We then evaluate benchmark and proposed methods—trained solely on these limited demonstrations—across arbitrarily translated and rotated test scenarios, thereby isolating and testing individual components of our spatially generalizable contact-rich policy.
Figure 2: (Left) Overview of the workspace for the peg-in-hole assembly task is presented. $2$ external cameras with calibrated extrinsics and $2$ wrists cameras are installed. (Right-Top) Peg and hole assembly with $1\mathrm{mm}$ of clearance. (Right-Bottom) Hole part with flat and tilted ($30^{\circ}$) platforms.
IV Solution Approach
We now describe our proposed EquiContact policy, which integrates a high-level vision-based planner—Diffusion Equivariant Descriptor Field (Diff-EDF)—with a low-level compliant visuomotor policy—Geometric Compliant control Action Chunking Transformer (G-CompACT). G-CompACT itself consists of a behavior cloning module built on a transformer-based CVAE architecture, followed by a geometric admittance controller (GAC). As discussed in Section III, because of space constraints, we focus on the insertion task in this paper, which will later be extended to include the picking task to complete the full pipeline implementation. We note that the Diff-EDF and G-CompACT are trained separately but are combined as a whole pipeline during inference time, as shown in Fig. 1.
IV-A Diffusion-Equivariant Descriptor Field (Diff-EDF)
Diff-EDF [14] is a diffusion-based manipulation policy with a bi-equivariant structure on $SE(3)$ transformations for pick-and-place tasks. Diff-EDF takes two point clouds as input: the scene point cloud $\mathcal{O}^{scene}$ captured by two external RGBD cameras with calibrated extrinsics (Fig. 2), and the gripper point cloud $\mathcal{O}^{grasp}$ in the end-effector frame, with or without a grasped object.
Diff-EDF outputs a pose $g_{{}_{EDF}}\in SE(3)$ representing the estimated target pose of the object of interest, which in our case is the hole position and is defined as:
$g_{{}_{EDF}}=f_{\theta_{1}}(\mathcal{O}^{scene},\mathcal{O}^{grasp})$ (1)
where the neural network $f_{\theta_{1}}$ implements the mapping. Due to the left-equivariant property, it follows that:
$g_{l}g_{{}_{EDF}}=f_{\theta_{1}}(g_{l}\circ\mathcal{O}^{scene},\mathcal{O}^{grasp}).$
Let $\mathcal{O}^{ref}\subset\mathcal{O}^{scene}$ denote the subset of points corresponding to the object of interest (e.g., the hole structure, see also Fig. 3). The model is designed to satisfy the following left-equivariance property on the point cloud characterization of the target object of interest:
Assumption 1 (Left-Equivariance of Diff-EDF).
$\begin{split}g_{{}_{EDF}}&=f_{\theta_{1}}(\mathcal{O}^{ref},\mathcal{O}^{grasp})\\ &\implies g_{l}\cdot g_{{}_{EDF}}=f_{\theta_{1}}(g_{l}\circ\mathcal{O}^{ref},\mathcal{O}^{grasp}),\end{split}$ (2)
To meet this assumption, we randomize the pose of the hole-platform assembly during training, both translationally and rotationally, with visual distractors in the background. This encourages the model to focus on $\mathcal{O}^{ref}$, not the rest of the objects in the scene.
While this property is not strictly enforced (due to the learned nature of $f_{\theta_{1}}$), Diff-EDF utilizes equivariant backbone architectures and localized attention mechanisms [31], enabling it to generalize from as few as $\sim 10$ demonstrations and have high robustness to left $SE(3)$ transformations.
To train Diff-EDF, the scene and grasp point clouds are collected together with the target reference frames, which represent the desired poses of the end-effector for pick-and-place operations. $15$ demonstrations were collected for the Diff-EDF: $12$ samples of the flat platform and $3$ samples of the tilted platform, both translationally and rotationally randomized. The training process of Diff-EDF follows the procedure in [14].
While the original Diff-EDF supports right-equivariance to handle transformations between the gripper and the grasped object, we did not utilize this feature. In practice, we observed that arbitrary peg transformations relative to the gripper introduced additional challenges. First, grasping the peg based on an arbitrary transformation leads to imprecise grasps, resulting in object slippage during contact. Second, the gripper-object rotation must be estimated continuously in real-time for G-CompACT to work under slippage. Instead, we enforced a consistent grasp orientation and relied only on the left-equivariance property. We refer the readers to [14, 10] for full details of Diff-EDF and to Supplementary Material  SI for implementation and heuristics. We emphasize again that Diff-EDF can be replaced by any point cloud-based method that outputs an approximate goal pose for (compliant) placing.
IV-B Geometric Compliant control Action Chunking with Transformers (G-CompACT)
IV-B1 Geometric Admittance Control (GAC)
We implement the geometric impedance control (GIC) proposed in [12, 13] in the geometric admittance control (GAC) setup [11]. We first note that we follow the notation used in [11]. Let the end-effector pose be denoted as $g\in SE(3)$ in a homogeneous matrix representation, or simply $g=(p,R)$, where $p\in\mathbb{R}^{3}$ is a position of the end-effector and $R\in SO(3)$ is a rotation matrix of the end-effector. Given the desired end-effector pose $g_{d}=(p_{d},R_{d})$, the desired end-effector dynamics for the GAC setup is given as follows:
$M\dot{V}^{b}+K_{d}V^{b}+f_{{}_{G}}=F_{e},$ (3)
where $M\in\mathbb{R}^{6\times 6}$ is symmetric positive definite desired inertia matrix, $K_{d}\in\mathbb{R}^{6\times 6}$ symmetric positive definite damping matrix, $F_{e}\in\mathbb{R}^{6}$ is external wrench applied to the end-effector in end-effector body frame and $V^{b}\in\mathbb{R}^{6}$ is a body-frame end-effector velocity. Further, $f_{{}_{G}}=f_{{}_{G}}(g,g_{d},K_{p},K_{R})\in\mathbb{R}^{6}$ is a elastic wrench given by:
$\begin{split}f_{{}_{G}}=\begin{bmatrix}f_{p}\\ f_{R}\end{bmatrix}=\begin{bmatrix}R^{T}R_{d}K_{p}R_{d}^{T}(p-p_{d})\\ (K_{R}R_{d}^{T}R-R^{T}R_{d}K_{R})^{\vee}\end{bmatrix},\end{split}$ (4)
where $\hat{f}_{{}_{G}}\in se^{*}(3)$, with $K_{p},K_{R}\in\mathbb{R}^{3\times 3}$ symmetric positive stiffness matrices for the translational and rotational dynamics, respectively, and $(\cdot)^{\vee}$ denotes the vee-map. By a slight abuse of notation we will frequently denote as $f_{{}_{G}}\in se^{*}(3)$ to emphasize that it is an elastic wrench on $SE(3)$, since $f_{{}_{G}}$ is isomorphic to $\hat{f}_{{}_{G}}$ [13].
Given the desired dynamics (3), the desired end-effector pose command $\tilde{g}_{d}(k)$ provided to the end-effector controller is calculated in discrete time as
$V_{d}^{b}(k)\!$ $=\!V^{b}(k)\!+\!T_{s}\cdot M^{-1}(F_{e}(k)\!-\!f_{{}_{G}}(k)\!-\!K_{d}V^{b}(k)),$
$\tilde{g}_{d}(k)\!$ $=\!g(k)\cdot\exp{(\hat{V}_{d}^{b}(k)\cdot T_{s})},$ (5)
where $T_{s}$ is a sampling time and $\hat{(\cdot)}$ denotes a hat-map. For the details of the GIC/GAC, we refer to [12, 11]. The admittance control loop is implemented at a $200\mathrm{Hz}$ rate using ROS2. We refer to Supplementary Material SI for detailed implementation tips.
IV-B2 G-CompACT
Action Chunking with Transformer (ACT) is a CVAE-based generative model designed for imitation learning in robotic manipulation tasks, originally introduced in [6]. While the original ACT formulation encodes robot proprioception using joint positions, this representation lacks spatial equivariance, limiting generalization. Instead, we adopt a proprioceptive encoding based on the end-effector pose, following the approach in [17].
In our previous work [11], we proposed a recipe for spatially equivariant policies, which involves a left invariance property and policy representation in the end-effector body frame. Following this principle and recent developments on relative action representations (similar to [29]), we structure our G-CompACT policy to take observations composed of: (1) Geometrically Consistent Error Vector (GCEV) proposed in [11], (2) FT sensor in the end-effector frame to capture contact behaviors, and (3) RGB images from wrist cameras. The model returns as action sequences: (1) relative pose from the current end-effector frame, and (2) admittance gains.
The GCEV $e_{{}_{G}}(g,g_{{}_{EDF}})$ is defined as
$e_{{}_{G}}(g,g_{{}_{EDF}})=\begin{bmatrix}R^{T}(p-p_{{}_{EDF}})\\ (R_{{}_{EDF}}^{T}R-R^{T}R_{{}_{EDF}})^{\vee}\end{bmatrix},$ (6)
where $g=(p,R)\in SE(3)$ is a current end-effector pose and $g_{{}_{EDF}}=(p_{{}_{EDF}},R_{{}_{EDF}})$ is the reference frame predicted by Diff-EDF. The derivation of the GCEV $e_{{}_{G}}$ is presented in [12]. We note that the GCEV denotes an error vector on the end-effector body frame as emphasized in [11].
The model outputs a chunk of $N$ relative pose actions:
$\bar{g}_{d}(k+i)=g(k)\cdot g_{rel}(k+i),\quad i=1,\cdots,N,$ (7)
where each $g_{rel}(k+i)\in SE(3)$ is the $i$th relative pose from the ACT model and $\bar{g}_{d}(k+i)$ is the $i$th temporary desired pose signal. A temporal ensembling mechanism is then applied to $\bar{g}_{d}(k+i)$ to compute the desired pose $g_{d}(k)$ for GAC. To respect the manifold structure of $SE(3)$, we use the scipy implementation for weighted averaging on the rotation matrix to get $R_{d}(k)$ and standard weighted averaging on positions for $p_{d}(k)$ separately.
Our proposed G-CompACT also outputs the admittance gains matrices $K_{p}$ and $K_{R}$ in the GAC elastic wrench in Eq. (4). Unlike CompACT, the gain matrices are represented in the desired frame of the end-effector [13], removing the necessity of utilizing full-gain matrix representations and coordinate transformations to deal with spatial equivariance. Therefore, a diagonal matrix representation in $K_{p}$ and $K_{R}$ is sufficient to conduct the PiH task successfully, without requiring a Cholesky factorization to represent the gains, as in CompACT. Since the diagonal component representation of matrices is equivalent to a vector, the standard temporal ensemble can be applied. Similar to the relative pose action, the ACT provides temporary gain signals $\bar{K}_{p}$ and $\bar{K}_{R}$, followed by a temporal ensemble to obtain $K_{p}$ and $K_{R}$.
Formally, let the G-CompACT method be $\pi_{\theta}$ such that
$\begin{split}a(k)&=\pi_{\theta_{2}}(o(k)),\quad\text{where}\\ a(k)&\triangleq\{(g_{rel}(k+i),\bar{K}_{p}(k+i),\bar{K}_{R}(k+i))\}_{i=1}^{N},\\ o(k)&\triangleq(e_{{}_{G}},F_{e},I_{w})(k),\end{split}$ (8)
where $\bar{K}_{p}$ and $\bar{K}_{R}$ denote the temporary admittance gains, $a(k)$ denotes the actions and $o(k)$ denotes the observation at time step $k$, $e_{{}_{G}}$ is GCEV defined in (6), $F_{e}$ is force-torque sensor reading on the end-effector frame, and $I_{w}=\{I_{w,1},I_{w,2}\}$ where $I_{w,1}$ and $I_{w,2}$ denote the images from the wrist cameras #1 and #2, respectively (Fig. 2).
Since the images $I_{w}$ are fed to the transformer decoder after being passed to the visual encoder structure, one can further consider the G-CompACT as
$\begin{split}a(k)&=\pi^{\prime}_{\theta_{3}}(e_{{}_{G}},F_{e},z)(k),\quad\text{where}\;z(k)=\mu_{\phi}(I_{w}).\end{split}$ (9)
where $\mu_{\phi}$ denotes the visual encoder, e.g., ResNet, and $z$ is the latent variable from the visual encoder.
We now introduce the following assumption regarding the left-invariant visual representation.
Assumption 2 (Left-invariant Visual Features).
The features from visual encoder $\mu_{\phi}$ are left invariant, i.e.,
$z(k)=\mu_{\phi}(g_{l}\circ I_{w})=\mu_{\phi}(I_{w}),\;\;\forall g_{l}\in SE(3).$ (10)
Note that we refer to $g_{l}$ as a left-group action [10]. The left-group action applied on the wrist-camera images is illustrated in Fig. 3. The meaning of the visual representation $z$ being left-invariant is that the vision encoder $\mu_{\phi}$ is trained to focus only on group action invariant features, such as the flat surface surrounding the hole on the platform. The satisfaction of Assumption 2 can be challenging. To satisfy this assumption, we have designed our platform and the surface surrounding the hole assembly so that there are sufficient surface features and the cameras primarily see the surface surrounding the hole, not the lower parts of the platform. This engineering choice was somewhat ad hoc, as there is no guarantee or sufficient inductive bias to encourage the desired behavior. We will demonstrate in later experimental results that this assumption may not apply in certain cases.
Figure 3: Effects of the left group action $g_{l}$ to the end-effector pose $g$ and the reference frame $g_{ref}$, and to the wrists cameras $I_{w,1}$ and $I_{w,2}$. As the left group action is applied on the end-effector, the wrist cameras start to see not only the optical table but also the outer backgrounds. $\{s\}$ denotes a spatial frame.
The following proposition shows the left-invariance of the G-CompACT method in the end-effector frame.
Proposition 1 (Left-invariance of G-CompACT).
Suppose that the Assumption 2 holds. Then,
$a(k)=\pi_{\theta_{2}}(g_{l}\circ o(k))=\pi_{\theta_{2}}(o(k)).$ (11)
Proof.
The left-translated observation signals $o(k)$ reads that:
$g_{l}\circ o(k)=(g_{l}\circ e_{{}_{G}},g_{l}\circ F_{e},g_{l}\circ I_{w}).$ (12)
As was shown in Lemma 1 of [11], the GCEV $e_{{}_{G}}$ is left invariant as
$g_{l}\circ e_{{}_{G}}(g,g_{{}_{EDF}})=e_{{}_{G}}(g_{l}g,g_{l}g_{{}_{EDF}})=e_{{}_{G}}(g,g_{{}_{EDF}}).$
The force-torque sensor values are left-invariant because they are already defined with respect to the end-effector frame [11], and the visual representation vectors satisfy left invariance due to Assumption 1. Combining all these properties, it follows that
$a(k)=\pi_{\theta_{2}}(g_{l}\circ o(k))=\pi^{\prime}_{\theta_{3}}(e_{{}_{G}},F_{e},z)=\pi_{\theta_{2}}(o(k)),$ (13)
which shows the left invariance of the G-CompACT policy on the end-effector frame. ∎
In what follows, we demonstrate that the pose signal produced by G-CompACT is equivariant when described in the spatial frame.
Corollary 1 ($SE(3)$ (Left) Equivariance of G-CompACT).
Let the output of G-CompACT in the spatial frame be $\bar{g}_{d}$ via (without considering the outputs related to gains)
$\bar{g}_{d}=g\!\cdot\!\pi_{\theta_{2}}(e_{{}_{G}}(g,g_{{}_{EDF}}),F_{e},I_{w}).$ (14)
Then,
$g_{l}\bar{g}_{d}$ $=g_{l}g\cdot\pi_{\theta_{2}}(g_{l}\circ e_{{}_{G}}(g,g_{{}_{EDF}}),g_{l}\circ F_{e},g_{l}\circ I_{w})$
$=g_{l}g\cdot\pi_{\theta_{2}}(e_{{}_{G}}(g,g_{{}_{EDF}}),F_{e},I_{w}).$ (15)
Since the temporary desired pose on the spatial frame signal, $\bar{g}_{d}$, is $SE(3)$ equivariant, the temporal aggregated signal $g_{d}$ is also equivariant. Therefore, by combining the G-CompACT with a spatial frame representation and temporal ensemble, we will now define the final form of G-CompACT as $\hat{\pi}_{\theta_{2}}$, which is given by:
$(g_{d},K_{p},K_{R})=\hat{\pi}_{\theta_{2}}(e_{{}_{G}}(g,g_{{}_{EDF}}),F_{e},I_{w}).$ (16)
and satisfies the following equivariance property:
$\begin{split}&(g_{l}g_{d},K_{p},K_{R})(k)=\hat{\pi}_{\theta_{2}}(g_{l}\circ o(k)).\end{split}$ (17)
To train a G-CompACT, we collect the expert demonstration using teleoperation on the fixed pose of the platform. During data collection, the expert teleoperator monitors the task’s progress, makes real-time movement commands via a SpaceMouse, and adjusts the admittance gains using keyboard input to switch between the following predefined gain modes: low-gain mode, high-gain mode, insertion mode, and contact mode. The low/high gain mode has low/high gains in all directions, the insertion mode has high gains in the $z$ direction of the end-effector frame and low gains elsewhere. Finally, the contact mode has low gains in the $z$ direction and high gains elsewhere. We collected $86$ demonstrations to train a near-perfect policy.
Since we know the fixed pose of the platform a priori, e.g., a ground-truth reference frame, the GCEV vector can be calculated for the training process. Nevertheless, the reference frame needs to be estimated via Diff-EDF (as $g_{{}_{EDF}}$) during the inference stage, which may have non-negligible errors. To handle this issue, we have added noise to the reference frame $g_{ref}$ to calculate $e_{{}_{G}}$ during dataset preprocessing. This provides the model with an inductive bias to primarily rely on $e_{{}_{G}}$ values for rough alignment and rely on vision feedback for fine-grained motion. The rest of the training follows the standard imitation learning pipeline.
IV-C EquiContact
The proposed EquiContact method comprises the high-level Diff-EDF, which serves as a high-level vision planner that provides the reference frame to be fed to GCEV, and the low-level G-CompACT, which handles fine-grained movement and contact interaction during insertion using real-time vision and force feedback. The overall pipeline of the EquiContact is presented in Fig. 1. Here, we demonstrate that the proposed EquiContact possesses the equivariance property. Let EquiContact be written as $h_{\theta}$ so that $h_{\theta}(g,g_{ref},F_{e})\mapsto f_{{}_{G}}$, i.e., $h_{\theta}:SE(3)\times SE(3)\times se^{*}(3)\to se^{*}(3)$.
Proposition 2.
The EquiContact Policy $h_{\theta}$ is equivariant if it is described relative to the spatial frame.
Proof.
Suppose that the Assumption 1 and 2 hold. Let the object of interest, e.g., a peg for the picking task and a hole for the placing task, be observed by $\mathcal{O}^{ref}$, $I_{w}$ with its pose given by $g_{ref}$, so that the left-translated $g_{l}\cdot g_{ref}$ is observed by $g_{l}\circ\mathcal{O}^{ref}$ from the point cloud, and $g_{l}\circ I_{w}$ by the left-translated end-effector attached wrist camera as described in Fig. 3.
First, notice that $h_{\theta}$ can be fully written as
$h_{\theta}(g,g_{ref},F_{e})$ $=f_{{}_{G}}(g,\hat{\pi}_{\theta_{2}}(e_{{}_{G}}(g,f_{\theta_{1}}(\mathcal{O}^{ref})),F_{e},I_{w}))$ (18)
$=f_{{}_{G}}(g,\underset{=(g_{d},K_{p},K_{R})}{\underbrace{\hat{\pi}_{\theta_{2}}(e_{{}_{G}}(g,g_{{}_{EDF}}),F_{e},I_{w})}}).$
Then, when both $g$ and $g_{ref}$ undergoes a left transformation $g_{l}$, from Assumption 1 and Corollary 1, the following holds:
$h_{\theta}(g_{l}g,g_{l}g_{ref},g_{l}\circ F_{e})$
$=f_{{}_{G}}(g_{l}g,\hat{\pi}_{\theta_{2}}(e_{{}_{G}}(g_{l}g,f_{\theta_{1}}(g_{l}\circ\mathcal{O}^{ref})),g_{l}\circ F_{e},g_{l}\circ I_{w}))$
$=f_{{}_{G}}(g_{l}g,\hat{\pi}_{\theta_{2}}(e_{{}_{G}}(g_{l}g,g_{l}g_{{}_{EDF}}),g_{l}\circ F_{e},g_{l}\circ I_{w}))$
$=f_{{}_{G}}(g_{l}g,g_{l}g_{d},K_{p},K_{R})=f_{{}_{G}}(g,g_{d},K_{p},K_{R})$
$=h_{\theta}(g,g_{ref},F_{e}).$ (19)
We note that the second-last equation ($SE(3)$ left-invariance of the elastic wrench) comes from Lemma 1 of [11]. Finally, from the result of Proposition 2 of [11], it follows that
$h^{s}_{\theta}(g_{l}g,g_{l}g_{ref},g_{l}\circ F_{e})=\text{Ad}_{g_{l}^{\text{-}1}}^{T}h_{\theta}^{s}(g,g_{ref},F_{e}),$ (20)
where Ad is an Adjoint operator, $\text{Ad}_{g_{l}^{\text{-}1}}^{T}$ is a representation for $se^{*}(3)$ (wrench) domain for the group action $g_{l}$ [11], and superscript $s$ denotes a vector represented in the spatial frame. ∎
IV-D Extensions to Pick Tasks
So far, we have described our method in terms of the insertion (placement) task. The proposed method can be extended to pick tasks in the same manner. The Diff-EDF can be utilized to obtain the pick reference frame, which is used for $e_{{}_{G}}$ for the picking G-CompACT. The picking G-CompACT is trained in such a way that the manipulator grasps a peg in a fixed, aligned pose, which helps EquiContact bypass the right-equivariance issue. For G-CompACT, the FT sensor values are not utilized as one of its observations, and it does not output the admittance gains; instead, it uses fixed gains.
V Experimental Results and Discussions
V-A Benchmark Results
To validate the effectiveness of our proposed EquiContact framework and its underlying design principles, we compare EquiContact against three benchmark approaches: ACT with world-frame observation and action representation, executed with and without GAC, and CompACT [17]. We note that the GAC serves as a standard admittance control when the platform’s pose is fixed, i.e., untilted. The observation and action representation of the benchmarks and the proposed EquiContact are listed in Table I. We conducted a series of benchmark tests to demonstrate the effectiveness of each component in the proposed approach. The overall results are summarized in Table I.
TABLE I: Success rates of the insertion policies in real-world experiments for the proposed and benchmark approaches. “In-Dist.” denotes in-distribution data and “OOD” denotes out-of-distribution data. For the In-Dist. (in distribution) scenario, the initial pose of the end-effector is randomized around the flat platform.
Methods Observation Action Cameras Test Scenario Success Rate
ACT w/o GAC [World Pose] [World Pose] 2 Wrists Cameras Flat Platform (In-Dist.) 2 / 10
ACT w/ GAC [World Pose] [World Pose] 2 Wrists Cameras Flat Platform (In-Dist.) 20 / 20
CompACT [World Pose, FT] [World Pose, Gains] 2 Wrists Cameras Flat Platform (In-Dist.) 20 / 20
Flat Platform (OOD) 0 / 10
EquiContact [GCEV, FT] [Relative Pose, Gains] 2 Wrists Cameras Flat Platform (OOD) 20 / 20
(Place, Ours) Tilted Platform ($30^{\circ}$, OOD) 19 / 20
V-A1 Demonstration of Compliance
We begin by evaluating the role of compliance using the same ACT model architecture, executed with and without the Geometric Admittance Control (GAC). Results for this comparison are shown in the 1st and 2nd rows of Table I. Without GAC, the ACT model shows significantly lower success rates. The failure mode of the ACT w/o GAC involves collision: as the robot approaches the platform, excessive contact forces trigger safety shutdowns, preventing the task from completing. Due to the safety issue, we limited the number of trials without GAC to 10. This result demonstrates that the presence of compliant control is almost a decision factor between success and failure for contact-rich tasks.
To further analyze the effects of compliant gain control, we compare ACT with GAC and CompACT under in-distribution (In-Dist.) platform settings. Although both methods achieve perfect success rates, their force profiles during insertion differ substantially. As shown in Fig. 4, CompACT, which outputs task-adaptive admittance gains based on force-torque feedback, consistently produces lower interaction forces in all directions. The effectiveness of the CompACT compared to the baseline ACT was already presented in [17].
Figure 4: Force profiles of CompACT and ACT with GAC (fixed gains) during insertion tasks are presented. The CompACT with force-torque sensor inputs and output gains shows lower interaction force in all directions.
V-A2 Demonstration of Equivariance
Although the CompACT succeeds in insertion tasks in trained scenarios without excessive force exertion, it fails to generalize to spatially unseen configurations. This is expected, as its observation and action representations are defined in the global spatial frame, which neither guarantees nor encourages equivariance. The result of applying CompACT to the translationally unseen cases is shown in the Table I $3$rd row - Flat Platform (OOD). Note that the flat platform is only utilized, meaning it is only randomized translationally. We tested for $10$ cases and did not conduct more tests because it resulted in a $0\%$ success rate.
In contrast, the proposed method (EquiContact) achieves perfect success rates on the translationally unseen flat platform, as can be seen in the $4$th row of Table I. As shown in Section IV, EquiContact has $SE(3)$ vision-to-force equivariance, achieving a near-perfect success rate, even on the tilted platform, which undergoes a full $SE(3)$ transformation. We attribute the single failure case to a large error from Diff-EDFs that exceeded the noise level applied during training.
TABLE II: Success Rates of the proposed EquiContact for a full pipeline of pick and place.
Test Scenario Success Rate Failure Cases
Flat Platform (OOD) 20 / 20 N/A
Tilted Platform ($30^{\circ}$, OOD) 18 / 20 1 Pick, 1 Place
The result of EquiContact for the full pick-and-place task is summarized in Table II. The EquiContact also demonstrates a near-perfect success rate in the full pick-and-place pipeline for peg-in-hole tasks. However, as the whole task is formulated in sequential stages, the error in the previous stage tends to propagate to the following stages, leading to slightly increased failure cases.
From these experimental results, we draw the following key takeaway.
Remark 1 (Anchoring a Localized Policy on the Reference Frame for Equivariance).
The G-CompACT policy $\pi_{\theta}$ operates exclusively on localized inputs: GCEV, FT values in the end-effector frame, and images from wrist cameras. Its outputs, relative poses, and admittance gains are likewise defined in the local end-effector frame. The localized policy is anchored to a reference frame generated by the Diff-EDF planner. This architectural design induces spatial generalization and $SE(3)$-equivariance. Under this principle, one may hypothesize the reason for the spatial generalizability of recent vision-language-action models, which utilize both wrist and external cameras. The wrist camera provides a localized policy, anchored in the reference frame obtained from the external cameras and proprioceptive information. In fact, learning a reference frame from proprioceptive information (e.g., joint position) and the external camera is subtle because of its black-box nature. In contrast, our method, which utilizes a point cloud-based reference frame approach, is more interpretable. Thus, we propose “anchoring localized policies on globally estimated reference frames” as a general framework for the divide-and-conquer philosophy to enhance spatial generalization and interpretability. In our case, this structure also enables provable $SE(3)$ equivariance via differential geometric design.
V-B Failure Cases and Performance Recovery via Data Augmentation
TABLE III: Success Rates of the G-CompACT (insertion) trained with the base dataset and augmented dataset for the flat platform under the presence of visual distractor scenario and tilted platform with large angle cases. The evaluation is conducted with the ground-truth reference frames.
Test Scenario Base Dataset Aug. Dataset
W/ Visual Distractor 4 / 10 9 / 10
Tilted Platform ($45^{\circ}$) 5 / 10 9 / 10
Although the EquiContact framework is vision-to-force equivariant in theory, this guarantee holds only under Assumptions 1 and 2. In particular, Assumption 2, which requires left-invariance of the visual features, is more challenging to enforce in practice, as our approach does not explicitly encode this property through loss functions or architectural inductive bias. As a result, the proposed G-CompACT algorithm shows degraded performance in scenarios with visual distractors or on a more severely tilted platform $(45^{\circ})$ – see Table III (“Base Dataset” column). Note that the larger tilting angle results in more out-of-distribution images, as the wrist cameras begin to capture more unseen background scenes, as illustrated in Fig. 3. To accommodate this, we augment the base dataset with $20$ demos collected using the visual distractor on a fixed, flat platform and $20$ demos collected from a $30^{\circ}$ angle of the tilted platform. After training with the augmented dataset, the success rates return to normal levels, as shown in the “Aug. Dataset” column of Table III, demonstrating the performance recovery achieved through data augmentation.
VI Conclusion
In this work, we introduced EquiContact, a vision-to-force equivariant policy for spatially generalizable contact-rich tasks. By integrating a global reference frame estimator (Diff-EDF) with a fully localized visuomotor servoing policy module (G-CompACT), we demonstrate how compliance, localized policy, and induced equivariance can be unified to enable the peg-in-hole (PiH) task, a representative contact-rich precision task, under spatial perturbations. We proved the $SE(3)$ equivariance of the policy under assumptions on point cloud and image observations and validated its effectiveness through real-world experiments on PiH benchmarks. Compared to benchmark methods, our approach generalizes to unseen platform positions and orientations while maintaining low contact force and near-perfect success rates. Through extensive benchmark studies, we highlighted the effectiveness of the three principles – compliance, localized policy, and induced equivariance – for achieving spatial generalizability in contact-rich manipulation. We conclude that these principles offer a simple yet powerful design guideline for developing spatially generalizable and interpretable robotic policies complementing recent trends in end-to-end visuomotor learning and enabling a structured divide-and-conquer approach.
Supplementary Material
SI Implementation Details
In this section, we provide the details of the hardware code implementation. Since every hardware system has a different implementation, our goal is to provide a comprehensive overview and a guide to our code distribution. We will use the exact names of the nodes to enable readers to easily understand our codebase.
SI-A Implementation Pipeline
The pipeline visualized in Fig. S1 is an interconnected system of camera, robot, and controller nodes with two policy services, all implemented in ROS2 Humble. The two policies are implemented as services because they are activated infrequently and latencies in activating the policies initially do not affect the behavior in practice. The execution of the full pipeline is divided into two primary stages: querying Diffusion-EDF for initial pick and insertion end-effector poses, and then executing the appropriate ACT sub-policies for these two steps after reaching a neighborhood of the predicted poses. In the first stage, a request to the Diffusion-EDF Service is made through the FullPipelineClient, which prompts the service to pull the individual point clouds published by the Orbbec Node. The service subsequently merges the point cloud in the spatial frame, preprocesses it, and passes it to the initialized Diff-EDF model to produce both pick and insertion poses. The Diff-EDF model, as covered in the following section, is formulated as a “double-pick” model that assumes that a fixed grip will be executed, and a cached empty gripper point cloud is passed along with the scene point cloud to get both predictions.
In the second stage, the FullPipelineClient requests the G-CompACT Service node to complete the task with these two poses as noisy references. The service initializes both pick and place policies and executes the tasks serially, raising flags to signal completion of each step. For raw camera images and proprioceptive information, the service subscribes to nodes for each camera (from Wrist Camera Nodes) and the Robot Node respectively; note that the controller and proprioceptive helper are encapsulated in one Robot Node that manages both publishing proprioceptive information and executing policy predictions. During the pick portion, the pick ACT model only publishes end-effector poses, and fixed high admittance gains are utilized because the demonstrations encourage the model to make limited contact with the peg. The pick model has a fixed time to complete the task before a grasp is attempted, and the robot moves to the next step. Similarly, the place ACT policy is executed given a fixed-time limit, but for the place model, the force-torque data is also passed as an input, and admittance gains are outputted along with the end-effector poses. A fixed-time limit for both phases is used because it is easily implemented and works more consistently than other, more complicated completion heuristics.
Figure S1: Whole pipeline implemented (via client named as FullPipelineClient) with ROS2 is presented.
SI-B Diffusion-EDF
SI-B1 Implementation Heuristics
Instead of following the original Diff-EDF pipeline, which utilizes pick-and-place models, we used two pick models. This decision mainly stems from the task setup, where the peg is upright, and the peg is grasped by the gripper in an aligned, upright pose. The core difference of the pick and place model is that the place model needs to get the grasp point cloud after each grasp to handle the right equivariance of the model. However, from task setup, we bypass this right equivariance issue, removing the necessity of the place model.
We also used the post-processing heuristics to filter the output pose of the Diff-EDF. The Diff-EDF outputs $20$ candidate target poses for picking and $10$ candidates for the place, which are ranked by the energy level. Although in theory, the lower energy poses should result in a better pose, we found out that this does not hold in practice. Instead, we figured out that the Diff-EDF have low variations on the position of the “tip”. To elaborate, suppose a picking policy returns the desired poses $g_{pick}^{i}=(p_{pick}^{i},R_{pick}^{i})$, which are the end-effector poses, not the tip, or Tool-Center-Point (TCP) poses. Then, the position of the tip estimated by Diff-EDF can be calculated via
$p_{tip,pick}^{i}=p_{pick}^{i}+R_{pick}^{i}p_{et,pick}$
where $p_{et,pick}\in\mathbb{R}^{3}$ denotes the coordinate of the gripper’s tip with respect to the end-effector body frame. Using this property and the task setup that the peg is always upright, which gives $R_{peg}$, we recalculate the Diff-EDF of the picking as
$p_{pick}=\bar{p}_{tip,pick}-R_{peg}^{T}p_{et,pick}$
where $\bar{p}_{tip,pick}$ is the mean of $p_{tip,pick}^{i}$ for $i=1,\cdots,N$. On the other hand, for placement, we did not make any assumptions about the orientation of the hole-platform structure. The heuristic we used for the place is as follows:
$p_{place}=\bar{p}_{tip,place}-\bar{R}_{place}^{T}p_{et,place},$
where $\bar{R}_{place}$ is the averaged rotation of $R_{place}^{i}$, and $\bar{p}_{tip,place}$ is the averaged position of the tip calculated similarly from the picking case.
SI-B2 Accuracy of the Diff-EDF
Here, we present detailed error values of the Diff-EDF evaluated on the training dataset. We note that we followed the implementation pipeline of the original Diff-EDF package, but just utilized the two pick models as described in Section SI-B1. The resulting root mean squared error (RMSE) values are shown in Table S1. The rotational errors are calculated via $R_{e}=R_{gt}R_{pick}^{T}$ and converted to $xyz$ euler angle representation, where $pick$ is replaced by $place$ for the placing task.
TABLE S1: RMSE error values of Diff-EDFs. The dimension of translational error is $\mathrm{mm}$ and the rotational error is $\mathrm{deg}$.
Pick Place
Translation $x$ ($\mathrm{mm}$) 16.76 4.981
Translation $y$ ($\mathrm{mm}$) 5.886 7.422
Translation $z$ ($\mathrm{mm}$) 10.79 5.236
Rotation $x$ ($\mathrm{deg}$) 19.88 13.74
Rotation $y$ ($\mathrm{deg}$) 40.55 18.99
Rotation $z$ ($\mathrm{deg}$) 106.7 91.41
As mentioned in Section I, the error in translation ranges from $5\mathrm{mm}$ to $15\mathrm{mm}$, and rotational error can range up to $40^{\circ}$. We note that the $z$-directional rotation error can be arbitrarily high in our case because our PiH task is symmetric.
SI-C Tips on applying Geometric Admittance Control
Here, we present a few tips on the application of geometric admittance control. The GAC is implemented as a part of a ROS2 node, running with $200\mathrm{Hz}$.
SI-C1 ROS2 Implementation
The Robot Node acts as the central controller, receiving and executing desired pose commands using a Geometric Admittance Control (GAC). Its main execution loop operates at $200\mathrm{Hz}$, ensuring smooth and compliant motion in response to external interactions. The node subscribes to the “action” topic, which contains the pose command and admittance gains command. Additionally, the controller subscribes to the “rebias” topic in order to preserve the accuracy of force readings. When a message on this topic is received, a callback raises a flag, triggering a re-biasing procedure for the FT sensor during the next control cycle. Within its main loop, the controller sends the intended pose and processed FT sensor data to the GAC function, which generates the compliant commands for the robot’s low-level motion controller.
SI-C2 Handling Force-Torque Sensor Data
The first consideration is how to handle the Force-Torque (FT) sensor attached to the end-effector. It is always favorable to subtract the bias of the FT sensor after finishing each task, such as picking up the peg. We collect the FT sensor value for $2$ seconds and take a mean value to use as a bias. The other is to filter the FT sensor. Since the FT sensor is noisy, it is strongly recommended to implement a low-pass filter. However, having a low cut-off frequency can lead to serious phase lag, resulting in so-called “bouncing” behavior. We designed a simple 1st-order low-pass filter and implemented it using a state-space representation in discrete time.
SI-C3 Implementation Tips
As mentioned in Section. IV, the detailed implementation method depends on the low-level servo controller type. Typical controllers provided by robot manufacturers are joint position control, joint velocity control, end-effector position control, and end-effector velocity control. If a torque controller is readily available, then directly implementing a geometric impedance controller (GIC) is recommended, but joint friction compensation needs to be handled appropriately. The end-effector velocity control can be straightforwardly applied from (IV-B1), by directly using $V_{d}^{b}(k)$ as a desired end-effector velocity command. One caveat here is the $V^{b}$, which is body-frame end-effector velocity. The common velocity provided by the manufacturer is defined on the Cartesian (Spatial) frame. If end-effector velocity is only offered in the Cartesian frame $V\in\mathbb{R}^{6}$, then
$V^{b}=\begin{bmatrix}R^{T}&0_{3\times 3}\\ 0_{3\times 3}&R^{T}\end{bmatrix}V,$
where $R$ is a rotation matrix from the current configuration. However, sometimes, it is provided in the Cartesian frame for translational velocity and the body frame for rotational velocity. The users should verify the frames of the velocities by running the robot in one direction and logging its values. For the joint space-based approach, one can use
$\begin{split}\dot{q}_{d}(k)\!&=\!\dot{q}(k)\!+\!T_{s}\cdot J_{b}(k)^{-1}M^{-1}(F_{e}(k)\!-\!f_{{}_{G}}(k)\!-\!K_{d}V^{b}(k)),\\ q_{d}(k)\!&=\!q(k)+T_{s}\cdot\dot{q}_{d}(k),\end{split}$
where $J_{b}$ denotes a body-frame Jacobian matrix. $\dot{q}_{d}(k)$ is provided to joint-velocity controller and $q_{d}(k)$ is provided to the joint-position controller. The Jacobian matrix also needs to be identified concerning the end-effector velocity, i.e., if the end-effector velocity is given in the Cartesian frame, then the Jacobian matrix $J$ is also provided in the Cartesian frame. In this case, the body-frame Jacobian matrix can be obtained via
$J_{b}=\begin{bmatrix}R^{T}&0_{3\times 3}\\ 0_{3\times 3}&R^{T}\end{bmatrix}J$
When the Jacobian matrix is not provided, an alternative is to use the robotics toolbox from Python or Matlab. First, use the DH parameters provided by the manufacturer (or URDF), then build the Jacobian matrix function. The function can be cast into a lambda function in Python. For Python, the Jacobian matrix function can also be wrapped into executable .so file using Matlab code builder and executed in the loop for fast computation (well below $1\mathrm{ms}$).
SI-D Teleoperation Details
The architecture of our teleoperation system is explained to make it easier to reproduce and adapt our work. The system is constructed modularly, with the human-operator interface decoupled from the robot controller. This allows greater adaptation to different robotic platforms and control algorithms. Two ROS nodes make up the majority of the architecture: the $\mathtt{RobotNode}$ (named as $\mathtt{robot\_gac}$ in the code implementation), which processes proprioceptive information and executes a Geometric Admittance Controller to carry out operations, and the $\mathtt{Teleoperation}$ (named as $\mathtt{teleop\_ros}$ in the code implementation) node, which handles operator input.
The teleoperation node acts as the principal interface between the human operator and the robot controller. A 6-DOF SpaceMouse is used to record input, which is then converted into the necessary end-effector pose. The main teleop function of this node operates at a frequency of $200\mathrm{Hz}$ inside a timer loop. It adjusts the desired end-effector pose incrementally during each cycle by reading velocity commands from the SpaceMouse.
The SpaceMouse provides velocity commands $V_{sm}=[v_{sm}^{T},\omega_{sm}^{T}]^{T}\in\mathbb{R}^{6}$, where $v_{sm}\in\mathbb{R}^{3}$ is translational velocity and $\omega_{sm}\in\mathbb{R}^{3}$ is rotational velocity. The change in desired position, $\Delta p$, is computed by scaling $v_{sm}$ with a position offset:
$\Delta p=v_{sm}\cdot\mathtt{pos\_offset}$
The target position is then updated as:
$p_{d}(k+1)=p_{d}(k)+\Delta p$
The desired end-effector orientation, $R_{d}(k+1)$, is updated by composing the previous rotation with the incremental rotation matrix, $\Delta R$. The small, incremental angles, $\Delta\theta$, are computed by scaling $w_{sm}$ with rotation offset:
$\Delta\theta=w_{sm}\cdot\mathtt{rot\_offset}$
The $\Delta\theta$ is then converted to $\Delta R$ using Euler angle as $\Delta R=\mathtt{euler2rotmat(\Delta\theta)}$. The target orientation is then updated as:
$R_{d}(k+1)=R_{d}(k)\cdot\Delta R$
This calculated pose is then published as a command for the robot controller at a frequency of $200\mathrm{Hz}$. To create a dataset of the teleoperation commands, a separate timer logs the same message at $30\mathrm{Hz}$. This implements a teleoperation that operates on the robot base frame for translational movement and the end-effector frame for rotational movement.
This teleoperation node was created with multiple operation modes to improve usability. For end-effector alignment with tilted platforms, the operator can switch the SpaceMouse’s control frame between the end-effector frame and the robot’s base frame. Additionally, a dedicated button allows users to modify the sensitivity offsets for position and rotation, enabling users to alternate between a translation-focused and an orientation-focused mode. The operator can globally scale motion sensitivity by cycling through low, medium, and high speed levels. The system also includes specific features, such as an “insertion mode” that operates in the end-effector frame for the translational movement to help with high-precision alignment tasks, as well as requests to set the robot’s end effector at “home” or a randomized pose.
Furthermore, the operator can dynamically adjust the parameters of the Geometric Admittance Controller via dedicated keyboard inputs. There are four gain profiles: Free mode applies high stiffness gains $(1000)$ to all axes for precise motion in open space; Contact mode uses low z-axis gain $(300)$ with high gain elsewhere $(1500)$ to exert less force on surfaces; Insertion mode reverses this, with high z-axis gain $(1500)$ for precision while allowing compliance in other directions $(300)$; and Compliant mode sets all gains low $(300)$ to make the end-effector highly responsive to external forces. These gain profiles are published as action messages, together with poses and gripper states.
SI-E G-CompACT Training Details
During the training of the G-CompACT, we added the noise to the known fixed pose of the hole and peg. As mentioned earlier, this is because the output of the Diff-EDFs $g_{{}_{EDF}}$ is biased. By adding the noise to the goal pose signal, we encourage the model to focus on the image of the wrist camera instead of the incorrect GCEV $e_{{}_{G}}$ signal. For our training, we added noise to every step of the demonstration dataset drawn from a uniform distribution. The noisy target poses $\tilde{g}_{ref}=(\tilde{p}_{ref},\tilde{R}_{ref})$are calculated as
$\begin{split}\tilde{p}_{ref}&=p_{ref}+n_{p},\\ \tilde{R}_{ref}&=R_{ref}\cdot\mathtt{euler2rotmat}(n_{r}),\end{split}$
where $n_{p}\sim\mathcal{U}(-0.02,0.02)\in\mathbb{R}^{3}$ and $n_{r}\sim\mathcal{U}(-8^{\circ},8^{\circ})\in\mathbb{R}^{3}$, which denotes $2\mathrm{cm}$ of translational and $8^{\circ}$ (degree) of rotational error.
To encode the element on $SE(3)$ for learning, we used position & rotation vector representation for G-CompACT. However, in benchmark models of CompACT and ACTs, we utilized position & rot6d representation. This is because the rotation vector representation has a sign flip issue when the rotation angle is near $180^{\circ}$, but the rot6d representation does not have that problem. For G-CompACT, as the actions are represented with respect to the current end-effector frame, the transformation angles are relatively small and do not have a sign flip problem.