### v2

# EquiContact: A Hierarchical SE(3) Vision-to-Force Equivariant Policy for Spatially Generalizable Contact-rich Tasks

Abstract
This paper presents a framework for learning vision-based robotic policies for contact-rich manipulation tasks that generalize spatially across task configurations. We focus on achieving robust spatial generalization of the policy for the peg-in-hole (PiH) task trained from a small number of demonstrations. We propose EquiContact, a hierarchical policy composed of a high-level vision planner (Diffusion Equivariant Descriptor Field, Diff-EDF) and a novel low-level compliant visuomotor policy (Geometric Compliant Action Chunking Transformers, G-CompACT). G-CompACT operates using only localized observations (geometrically consistent error vectors (GCEV), force-torque readings, and wrist-mounted RGB images) and produces actions defined in the end-effector frame. Through these design choices, we show that the entire EquiContact pipeline is $SE(3)$-equivariant, from perception to force control. We also outline three key components for spatially generalizable contact-rich policies: compliance, localized policies, and induced equivariance. Real-world experiments on PiH, screwing, and surface wiping tasks demonstrate a near-perfect success rate and robust generalization to unseen spatial configurations, validating the proposed framework and principles. The experimental videos can be found on the project website: https://sites.google.com/berkeley.edu/equicontact
I Introduction
Imitation learning has recently shown significant success in expanding the capabilities of machine learning in real-world robotics applications [1, 2]. In the early stages of robot learning, many methods formulated manipulation as sequences of keyframe-based pick-and-place actions [3, 4]. More works have started to produce a continuous set of actions directly from vision inputs [5, 6, 7]. Similar to the trend seen in large language models (LLMs), there is a growing belief that large-scale data can unlock generalizable, vision-based policies for robotics [8]. This has led to massive efforts to build large datasets [1] for training policies with general knowledge.
However, such policies often lack spatial generalizability and therefore require a large amount of data to learn robust behaviors. As described in [9], both action chunking transformers (ACT) [6] and diffusion policy (DP) [5] are evaluated only within the limited spatial variations. Furthermore, both methods exhibit near-linear performance growth as the demonstration dataset size increases, suggesting that the trained policies do not inherently generalize well to new spatial configurations, but rather tend to interpolate between seen demonstrations.
An alternative line of recent research focuses on leveraging symmetry—particularly equivariance—to enhance spatial generalizability, thereby improving sample efficiency during training [10, 11]. This approach requires fewer data but comes with its own challenges. Equivariant neural networks, being of a more specialized nature, are often not as well-developed and are more computationally intensive than their non-equivariant counterparts, making real-time and large-scale deployment more difficult. As a result, it becomes more attractive for users to use standard models trained with massive datasets in many instances.
In [11], a $SE(3)$-equivariant gain-scheduling policy using geometric impedance control (GIC) [12, 13] was proposed to solve peg-in-hole (PiH) problems. Inspired by the view that many manipulation tasks can be framed as pick-and-place problems [4], we modeled PiH as a compliant pick-and-place task, where final peg poses are provided by vision-based $SE(3)$-equivariant models such as Diffusion-EDF (Diff-EDF) [14]. Since both the high-level planner and low-level variable impedance controller are equivariant, they can be combined to form a vision-to-force equivariant policy. However, in practice, Diff-EDF’s placement accuracy proved insufficient for tight insertion tasks, which require sub-millimeter precision (details in Appendix -A). This revealed a key limitation: high-level vision planners may capture global structures but struggle with precision and contact-sensitive execution. From this insight, we introduce an intermediate layer between the planner and the low-level controller, which provides real-time visual feedback to correct the residual errors of the high-level planner.
In this paper, we propose EquiContact, a hierarchical $SE(3)$ vision-to-force equivariant policy for spatially generalizable, contact-rich tasks. It consists of two main components: a high-level planner using Diffusion Equivariant Descriptor Fields (Diff-EDF) [14], which estimates a local reference frame from point clouds, and a low-level compliant visuomotor policy based on Action Chunking Transformer (ACT) [6], which we refer to as Geometric Compliant ACT (G-CompACT). A key design feature of G-CompACT is that it only relies on local information: the force-torque signal in the end-effector frame, a geometrically consistent error vector (GCEV) [11], and wrist camera inputs. The output of G-CompACT is the relative desired pose and admittance gains, which are then sent to the geometric admittance controller (GAC) module to execute compliant control. Our contribution lies in the framework design, not in specific model choices; for example, Diff-EDF could be replaced by ET-SEED [15], or ACT by other visuomotor policies.
The main contributions of this paper are as follows:
  1. 1.
    We propose EquiContact, a hierarchical, provably $SE(3)$-equivariant policy from point clouds and RGB inputs to interaction forces for executing contact-rich tasks.
  2. 2.
    We identify three key principles for spatially generalizable contact-rich manipulation: left-invariant compliant control action (via GAC [11]), localized policy (left invariance), and induced equivariance. These enable $SE(3)$-equivariant behavior without requiring explicitly equivariant neural networks [10].
  3. 3.
    Under these principles, we present the necessary conditions for $SE(3)$ vision-to-force equivariant policy, and mathematically prove the equivariance property of EquiContact.
  4. 4.
    We demonstrate that EquiContact achieves near-perfect success rates and spatial generalizability when these conditions are met in real robot experiments involving peg-in-hole, screwing, and surface wiping tasks.
From these key principles, we propose a general framework to enhance the spatial generalization and interpretability of vision-based policies, namely, “anchoring localized policy on globally estimated reference frame.” We emphasize that our work provides complementary insights to recent trends of robot learning [1, 8, 16, 2] that aim to build generalist policies from a large-scale demo dataset. Our principles offer structural guidelines to improve spatial generalizability through $SE(3)$ equivariance.
II Related Works
Visuomotor Servoing Methods Recently, generative modeling has become mainstream in realizing visuomotor servoing policies. Particularly, there are two dominant methods for visuomotor servoing: Action Chunking with Transformers (ACT) [6] and Diffusion Policy (DP) [5]. ACT utilizes the conditional variational autoencoder (CVAE) method as a generative model, whereas DP employs denoising diffusion as its generative model. ACT and DP have been extended to other approaches, including compliance and force-reactive behaviors [17, 18, 19], as well as structural improvements [20, 16, 21]. Our work is most closely related to CompliantACT (CompACT) [17], which integrates compliant control for visuomotor policies. We have significantly improved CompACT by incorporating a provable $SE(3)$ equivariant structure.
Equivariant Methods Earlier equivariant approaches attempted to handle manipulation tasks as an extension of pick-and-place tasks, by leveraging $SE(3)$ equivariance from point clouds [22, 15, 14] or $SO(2)$ equivariance [3] from top-down views. Equivariant approaches have been extended to visuomotor policies, such as DP or flow matching, [23, 24], using point clouds. [25, 26] proposed $SO(2)$ equivariant visuomotor policies using 2D images, not fully considering $SE(3)$. In contrast, our approach induces full $SE(3)$ equivariance from vision to control force without relying on explicitly equivariant neural networks, but using structured observations and actions via geometrical canonicalization. Furthermore, by integrating with $SE(3)$ equivariant control, we generalize beyond table-top settings to contact-rich manipulation.
Manipulation in Object Frame Our induced $SE(3)$ equivariant approach relies on representing the visuomotor policy in the end-effector frame. While recent works [27, 28, 29] define policies in the target (object) frame, policy representation in the end-effector frame offers improved fidelity and robustness. This is because the estimated object frame can be noisy, and the end-effector frame is reliably obtained via forward kinematics. Importantly, compared to [27, 28], we explicitly link the choice of reference frame to the equivariance property, and unlike [29], which only handles translational transformations, our method can deal with full $SE(3)$ transformations of the reference frame.
Figure 1: We propose an EquiContact, a hierarchical, provably $SE(3)$ vision-to-force equivariant policy for spatially generalizable contact-rich tasks. The proposed EquiContact consists of G-CompACT (Section. IV-B) and Diffusion-EDF (Section. IV-A). The G-Compact plays a localized policy over the reference frame provided by the Diffusion-EDF, making our framework generalizable to unseen task transformations during evaluation. The overall EquiContact algorithm is summarized in Algorithm 1.
III Problem Definition
In this paper, we aim to identify the key structural components required for learning policies that generalize spatially in contact-rich manipulation tasks. We will first focus on the peg-in-hole (PiH) problem as a representative force-based assembly task and validate the feasibility of the proposed approach to other contact-rich tasks later. Our proposed framework achieves $SE(3)$ vision-to-force equivariance through three essential design principles: (1) left-invariant compliant control, (2) localized policy, and (3) induced equivariance. These principles are validated through specific data collection and evaluation setups, as detailed below.
Unlike prior work [11] that assumes a known hole pose and a pre-grasped peg, we consider a more general setup: the robot must first grasp the peg and then perform insertion using only vision and proprioception, as illustrated in Fig. 2. We assume only that the peg is upright in the scene. Inspired by the idea that many manipulation tasks can be viewed as sequences of pick-and-place actions, we formulate insertion as a compliant pick-and-place task. However, high-level vision planners often lack the precision needed for tight-tolerance tasks like PiH, which in our case requires $1\mathrm{mm}$ accuracy.
To accommodate this, we propose a low-level compliant policy that:
  • provides real-time visual feedback to refine the coarse high-level command,
  • handles fine force-based interactions through compliance,
  • and achieves provable $SE(3)$-equivariance.
We assume that a high-level vision planner (e.g., Diff-EDF) can generate approximate pick-and-place poses. Our focus is on developing an equivariant compliant placing policy, i.e., insertion policy, using imitation learning. We further assume the peg is approximately aligned with the gripper during placement, since arbitrary peg poses introduce two challenges: (1) imprecise grasps can lead to slippage during contact, and (2) compensating for slippage requires continuous estimation of the gripper-to-peg transformation, which is difficult to achieve reliably in real time.
To train this policy, we collect expert demonstrations of insertions on a fixed-platform setting with a known hole location, where its objective is to train a policy that performs nearly perfectly in the trained scenario. We then evaluate benchmark and proposed methods—trained solely on these limited demonstrations—across arbitrarily translated and rotated test scenarios, thereby isolating and testing individual components of our spatially generalizable contact-rich policy.
Figure 2: (Left) Overview of the workspace for the peg-in-hole assembly task is presented. 22 external cameras with calibrated extrinsics and 22 wrists cameras are installed. (Right-Top) Peg and hole assembly with $1\mathrm{mm}$ of clearance. (Right-Bottom) Hole part with flat and tilted ($30^{\circ}$) platforms.
IV Solution Approach
We now describe our proposed EquiContact policy, which integrates a high-level vision planner—Diffusion Equivariant Descriptor Field (Diff-EDF)—with a low-level compliant visuomotor policy—Geometric Compliant ACT (G-CompACT). G-CompACT consists of a behavior cloning module based on a transformer-based CVAE architecture, followed by a geometric admittance controller (GAC). As discussed in Section III, we focus on the insertion task, with extension to picking addressed later in the paper. The Diff-EDF and G-CompACT are trained separately and combined at inference, as shown in Fig. 1.
Conceptually, EquiContact follows a simple yet powerful principle: “anchoring a localized policy on a globally estimated reference frame.” In our framework, G-CompACT serves as the localized low-level policy, operating solely on observations defined in the current end-effector frame; wrist camera RGB images, force-torque (FT) signals, and a geometrically consistent error vector (GCEV) $e_{{}_{G}}$. It outputs the desired relative pose and admittance gains, also in the end-effector frame. As both inputs and outputs are defined locally, G-CompACT functions as a fully localized policy. The high-level planner, Diff-EDF, estimates the pose of the target (e.g., hole) in the world frame. At the inference, the robot moves near the estimated reference frame, and G-CompACT is activated to perform compliant motion using only local feedback. Because the low-level policy does not depend on absolute global inputs, it can transfer robustly to unseen spatial configurations when the estimated reference frame is provided. This divide-and-conquer design provides a general framework for enhancing both spatial generalization and policy interpretability for contact-rich, and more broadly, general manipulation tasks.
In the remainder of this section, we formalize this spatial generalization property as $SE(3)$ equivariance and show how EquiContact satisfies this by construction through a rigorous mathematical derivation. We note that $\theta$, $\theta_{i}$, and $\phi$ with $i=1,2,3$ denote a learnable parameter of the neural network, without loss of generality. We also note that we follow the notation and definition conventions of [11, 30].
IV-A Diffusion-Equivariant Descriptor Field (Diff-EDF)
Diff-EDF [14] is a diffusion-based manipulation policy with a bi-equivariant structure on $SE(3)$ transformations for pick-and-place tasks. Diff-EDF takes two point clouds as input: the scene point cloud $\mathcal{O}^{scene}$ captured by two external RGBD cameras with calibrated extrinsics (Fig. 2), and the gripper point cloud $\mathcal{O}^{grasp}$ in the end-effector frame, with or without a grasped object.
Diff-EDF outputs a pose $g_{{}_{EDF}}\in SE(3)$ representing the estimated target pose of the object of interest, which in our case is the hole position and is defined as:
$g_{{}_{EDF}}=f_{\theta_{1}}(\mathcal{O}^{scene},\mathcal{O}^{grasp})$ (1)
where the neural network $f_{\theta_{1}}$ implements the mapping. Due to the left-equivariant property, it follows that: $g_{l}g_{{}_{EDF}}=f_{\theta_{1}}(g_{l}\circ\mathcal{O}^{scene},\mathcal{O}^{grasp}).$
Let $\mathcal{O}^{ref}\subset\mathcal{O}^{scene}$ denote the subset of points corresponding to the object of interest (e.g., the hole structure, see also Fig. 3). The model is designed to satisfy the following left-equivariance property on the point cloud characterization of the target object of interest:
Assumption 1 (Left-Equivariance of Diff-EDF).
$\begin{split}g_{{}_{EDF}}&=f_{\theta_{1}}(\mathcal{O}^{ref},\mathcal{O}^{grasp})\\ &\implies g_{l}\cdot g_{{}_{EDF}}=f_{\theta_{1}}(g_{l}\circ\mathcal{O}^{ref},\mathcal{O}^{grasp}),\end{split}$ (2)
To meet this assumption, we randomize the pose of the hole-platform assembly during training, both translationally and rotationally, with visual distractors in the background. This encourages the model to focus on $\mathcal{O}^{ref}$, not the rest of the objects in the scene. While this property is not strictly enforced (due to the learned nature of $f_{\theta_{1}}$), Diff-EDF utilizes equivariant backbone architectures and localized attention mechanisms [31], enabling it to generalize from as few as $\sim 10$ demonstrations and has high robustness to left $SE(3)$ transformations.
To train Diff-EDF, the scene and grasp point clouds are collected together with the target reference frames, which represent the desired poses of the end-effector for pick-and-place operations. $15$ demonstrations were collected for the Diff-EDF: $12$ samples of the flat platform and $3$ samples of the tilted platform, both translationally and rotationally randomized. The training process of Diff-EDF follows the procedure in [14].
Note that arbitrary peg transformations relative to the gripper can be handled via the right-equivariance feature of the Diff-EDF; however, due to the complications mentioned in Section. III, we enforced a consistent grasp and only utilized the left-invariance property. We refer the readers to [14, 10] for full details of Diff-EDF.
IV-B Geometric Compliant control Action Chunking with Transformers (G-CompACT)
IV-B1 G-CompACT
G-CompACT is based on the Action Chunking with Transformer (ACT), which is a CVAE-based generative model designed for imitation learning in robotic manipulation tasks [6]. To make G-CompACT spatially equivariant, we follow the principles proposed in [11]: left-invariant policy and policy representation in the end-effector body frame. To achieve this, we structure the observations and actions of the G-CompACT as described below.
First, the observations are given by, (1) Geometrically Consistent Error Vector (GCEV) $e_{{}_{G}}$ proposed in [11], (2) FT sensor in the end-effector frame $F_{e}$ to capture contact behaviors, and (3) RGB images $I_{w}=\{I_{w,1},I_{w,2}\}$ from wrist cameras (see Fig. 3). For actions, we choose (1) relative pose from the current end-effector frame $g_{rel}$, and (2) admittance gains for Geometric Admittance Control (GAC) $(K_{p},K_{R})$. The details of GAC and the definition of gains will be provided later in the Section.
Formally, let the G-CompACT method be $\pi_{\theta_{2}}$ such that
$\begin{split}a(k)&=\pi_{\theta_{2}}(o(k)),\quad\text{where}\\ a(k)&\triangleq(g_{rel},K_{p},K_{R})(k),\quad o(k)\triangleq(e_{{}_{G}},F_{e},I_{w})(k),\end{split}$ (3)
where $a(k)$ denotes the actions, and $o(k)$ denotes the observation at time step $k$. In fact, our G-CompACT also outputs the actions of chunk size $N$ following the standard action-chunking strategy. However, for the compact presentation, we will only consider the single-step action after undergoing proper processing, such as a temporal ensemble. The GCEV $e_{{}_{G}}(g,g_{{}_{EDF}})$ is defined as
$e_{{}_{G}}(g,g_{{}_{EDF}})=\begin{bmatrix}R^{T}(p-p_{{}_{EDF}})\\ (R_{{}_{EDF}}^{T}R-R^{T}R_{{}_{EDF}})^{\vee}\end{bmatrix}$, (4)
where $g=(p,R)\in SE(3)$ is a current end-effector pose, $g_{{}_{EDF}}=(p_{{}_{EDF}},R_{{}_{EDF}})$ is the reference frame predicted by Diff-EDF, and $(\cdot)^{\vee}$ denotes the vee-map, a mapping from $so(3)$ (Lie algebra of $SO(3)$) to $\mathbb{R}^{3}$. The physical meaning of GCEV is an error vector in the current end-effector frame between the current end-effector pose and the reference frame, considering the $SE(3)$ manifold structure, i.e., non-Euclidean. For the details of GCEV $e_{{}_{G}}$, we refer to [11, 12].
Since the images $I_{w}$ are fed to the transformer decoder after being passed to the visual encoder structure, one can further consider the G-CompACT as
$\begin{split}a(k)&=\pi^{\prime}_{\theta_{3}}(e_{{}_{G}},F_{e},z)(k),\quad\text{where}\;z(k)=\mu_{\phi}(I_{w}).\end{split}$ (5)
where $\mu_{\phi}$ denotes the visual encoder, e.g., ResNet, and $z$ is the latent variable from the visual encoder.
We now introduce the following assumption regarding the left-invariant visual representation.
Assumption 2 (Left-invariant Visual Features).
The features from visual encoder $\mu_{\phi}$ are left invariant, i.e.,
$z(k)=\mu_{\phi}(g_{l}\circ I_{w})=\mu_{\phi}(I_{w}),\;\;\forall g_{l}\in SE(3).$ (6)
Note that we refer to $g_{l}$ as a left-group action [10]. The left-group action applied on the wrist-camera images is illustrated in Fig. 3. The meaning of the visual representation $z$ being left-invariant is that the vision encoder $\mu_{\phi}$ is trained to focus only on group action invariant features, such as the flat surface surrounding the hole on the platform. The satisfaction of Assumption 2 can be challenging. To satisfy this assumption, we have designed our platform and the surface surrounding the hole assembly so that there are sufficient surface features and the cameras primarily see the surface surrounding the hole, not the lower parts of the platform. This engineering choice was somewhat ad hoc, as there is no guarantee or sufficient inductive bias to encourage the desired behavior. We will demonstrate in later experimental results that this assumption may not apply in certain cases.
Figure 3: Effects of the left group action $g_{l}$ to the end-effector pose $g$ and the reference frame $g_{ref}$, and to the wrists cameras $I_{w,1}$ and $I_{w,2}$. As the left group action is applied on the end-effector, the wrist cameras start to see backgrounds other than the optical table. $\{s\}$ denotes a spatial frame, i.e., robot base frame or world frame.
The following proposition shows the left-invariance of the G-CompACT method in the end-effector frame.
Proposition 1 (Left-invariance of G-CompACT).
Suppose that the Assumption 2 holds. Then,
$a(k)=\pi_{\theta_{2}}(g_{l}\circ o(k))=\pi_{\theta_{2}}(o(k)).$ (7)
The proof is presented in Appendix. -B. In what follows, we present an equivariant property of the pose signal produced by G-CompACT when described in the spatial frame.
Corollary 1 ($SE(3)$ Equivariance of G-CompACT).
The G-CompACT $\pi_{\theta_{2}}$ represented in the spatial frame satisfies the following equivariance property:
$\begin{split}&(g_{l}g_{d},K_{p},K_{R})(k)=\pi_{\theta_{2}}(g_{l}\circ o(k)).\end{split}$ (8)
Proof.
Notice that the desired pose signal in the spatial frame $g_{d}$ is obtained via $g_{d}=gg_{rel}=g\pi_{\theta_{2}}(o(k))$. Then, utilizing the left-invariance property, $g_{l}g\cdot\pi_{\theta_{2}}(g_{l}\circ o(k))=g_{l}g_{d}$. ∎
To train a G-CompACT, we collect the expert demonstration using teleoperation on the fixed pose of the platform. During data collection, the expert teleoperator monitors the task’s progress, makes real-time movement commands via a SpaceMouse, and adjusts the admittance gains using keyboard input to switch between the following predefined gain modes: low-gain mode, high-gain mode, insertion mode, and contact mode. The low/high gain mode has low/high gains in all directions, the insertion mode has high gains in the $z$ direction of the end-effector frame and low gains elsewhere. Finally, the contact mode has low gains in the $z$ direction and high gains elsewhere. We collected $86$ demonstrations to train a near-perfect policy.
Since we know the fixed pose of the platform a priori, e.g., a ground-truth reference frame, the GCEV vector can be calculated for the training process. Nevertheless, the reference frame needs to be estimated via Diff-EDF (as $g_{{}_{EDF}}$) during the inference stage, which may have non-negligible errors. To handle this issue, we have added noise to the reference frame $g_{ref}$ to calculate $e_{{}_{G}}$ during dataset preprocessing. This provides the model with an inductive bias to primarily rely on $e_{{}_{G}}$ values for rough alignment and rely on vision feedback for fine-grained motion. The rest of the training follows the standard imitation learning pipeline.
IV-B2 Geometric Admittance Control (GAC)
We implement the geometric impedance control (GIC) proposed in [12, 13] in the geometric admittance control (GAC) setup [11]. Let the end-effector pose be denoted as $g\in SE(3)$ in a homogeneous matrix representation, or simply $g=(p,R)$, where $p\in\mathbb{R}^{3}$ is a position of the end-effector and $R\in SO(3)$ is a rotation matrix of the end-effector. The GAC operates with the $(g_{d},K_{p},K_{R})$ signal calculated from G-CompACT, where the desired end-effector pose is calculated via $g_{d}=gg_{rel}$. Given $g_{d}=(p_{d},R_{d})$, the desired end-effector dynamics for the GAC setup is written as follows:
$M\dot{V}^{b}+K_{d}V^{b}+f_{{}_{G}}=F_{e},$ (9)
where $M\in\mathbb{R}^{6\times 6}$ is symmetric positive definite desired inertia matrix, $K_{d}\in\mathbb{R}^{6\times 6}$ symmetric positive definite damping matrix, $F_{e}\in\mathbb{R}^{6}$ is external wrench applied to the end-effector in end-effector body frame and $V^{b}\in\mathbb{R}^{6}$ is a body-frame end-effector velocity. $K_{d}$ matrix is selected to ensure overdamped system, as $K_{d}=3\cdot\text{blkdiag}(\sqrt{K_{p}},\sqrt{K_{R}})$. Further, $f_{{}_{G}}=f_{{}_{G}}(g,g_{d},K_{p},K_{R})\in\mathbb{R}^{6}$ is a elastic wrench given by:
[…]
where $K_{p},K_{R}\in\mathbb{R}^{3\times 3}$ symmetric being positive stiffness matrices for the translational and rotational dynamics, respectively. Following the standard admittance control formulation, the desired end-effector pose command is calculated using (9), which is passed to the lowest level tracking controller. For details on GIC/GAC, we refer readers to [12, 11]. The admittance control loop is implemented at a $200\mathrm{Hz}$ rate using ROS2.
IV-C EquiContact
The proposed EquiContact method comprises the high-level Diff-EDF, which serves as a high-level vision planner that provides the reference frame to be fed to GCEV, and the low-level G-CompACT, which handles fine-grained movement and contact interaction during insertion using real-time vision and force feedback. The overall pipeline of the EquiContact is presented in Fig. 1, and also summarized in Algorithm. 1. In Proposition 2, we demonstrate that the proposed EquiContact possesses the equivariance property. Let EquiContact be written as $h_{\theta}$ so that $h_{\theta}(g,g_{ref},F_{e})\mapsto f_{{}_{G}}$, i.e., $h_{\theta}:SE(3)\times SE(3)\times\mathbb{R}^{6}\to\mathbb{R}^{6}$.
Proposition 2.
Suppose that the Assumption 1 and 2 hold. The EquiContact Policy $h_{\theta}$ is equivariant if it is described relative to the spatial frame.
The proof is shown in the Appendix. -B.
Algorithm 1 Inference Procedure of EquiContact
1:Diff-EDF $f_{\theta_{1}}$, G-CompACT $\pi_{\theta_{2}}$, Task $\in$ {pick, place}
2:Get scene and grasp point cloud $\mathcal{O}^{scene}$, $\mathcal{O}^{grasp}$
3:Run Diff-EDF for reference frame $g_{{}_{EDF}}\!=\!f_{\theta_{1}}(\mathcal{O}^{scene},\mathcal{O}^{grasp})$
4:Move the end-effector near the reference frame and
5:   initialize EquiContact $\pi_{\theta_{2}}$
6:for each inference timestep $k$ do
7:  Get current sensor values $g(k)$, $F_{e}(k)$, $I_{w}(k)$
8:  Calculate GCEV $e_{{}_{G}}(k)=e_{{}_{G}}(g(k),g_{{}_{EDF}})$ (4)
9:  Run G-CompACT:
10:      $(g_{rel},K_{p},K_{R})(k)=\pi_{\theta_{2}}(e_{{}_{G}},F_{e},I_{w})(k)$ (3)
11:  Calculate desired pose $g_{d}(k)=g(k)g_{rel}(k)$
12:  Update $(g_{d},K_{p},K_{R})(k)$ for GAC loop
13:  Run GAC realizing desired dynamics (9), (10)
14:end for
IV-D Extensions to Pick Tasks
So far, we have described our method in terms of the insertion (placement) task. The proposed method can be extended to pick tasks in the same manner. The Diff-EDF can be utilized to obtain the pick reference frame, which is used for $e_{{}_{G}}$ for the picking G-CompACT. The picking G-CompACT is trained in such a way that the manipulator grasps a peg in a fixed, aligned pose, which helps EquiContact bypass the right-equivariance issue. For G-CompACT, the FT sensor values are not utilized as one of its observations, and it does not output the admittance gains; instead, it uses fixed gains.
V Experimental Results and Discussions
V-A Benchmark Results
To evaluate the effectiveness of EquiContact and its core design principles, we compare it against three baselines: ACT with world-frame observations and actions, executed with and without GAC, and CompACT [17]. Note that GAC reduces to standard admittance control when the platform is fixed, i.e., not tilted. Table. I summarizes the observation/action representations used in each method and reports the benchmark results across all setups.
TABLE I: Success rates of the insertion policies in real-world experiments for the proposed and benchmark approaches. “In-Dist.” denotes in-distribution data and “OOD” denotes out-of-distribution data. For the In-Dist. (in distribution) scenario, the initial pose of the end-effector is randomized around the flat platform.
Methods Observation Action Cameras Test Scenario Success Rate
ACT w/o GAC [World Pose] [World Pose] 2 Wrists Cameras Flat Platform (In-Dist.) 2 / 10
ACT w/ GAC [World Pose] [World Pose] 2 Wrists Cameras Flat Platform (In-Dist.) 20 / 20
CompACT [World Pose, FT] [World Pose, Gains] 2 Wrists Cameras Flat Platform (In-Dist.) 20 / 20
Flat Platform (OOD) 0 / 10
EquiContact [GCEV, FT] [Relative Pose, Gains] 2 Wrists Cameras Flat Platform (OOD) 20 / 20
(Place, Ours) Tilted Platform ($30^{\circ}$, OOD) 19 / 20
V-A1 Demonstration of Compliance
As the importance of left-invariant compliant control action has already been verified in [11], we focus on verifying the necessity of compliant control action. We begin by evaluating the role of compliance using the same ACT model architecture, executed with and without the Geometric Admittance Control (GAC). Results for this comparison are shown in the 1st and 2nd rows of Table I. Without GAC, the ACT model shows significantly lower success rates. The failure mode of the ACT w/o GAC involves collision: as the robot approaches the platform, excessive contact forces trigger safety shutdowns, preventing the task from completing. Due to the safety issue, we limited the number of trials without GAC to 10. This result demonstrates that compliant control is nearly a deciding factor between success and failure in contact-rich tasks.
To further analyze the effects of compliant gain control, we compare ACT with GAC and CompACT under in-distribution (In-Dist.) platform settings. Although both methods achieve perfect success rates, their force profiles during insertion differ substantially. As shown in Fig. 4, CompACT, which outputs task-adaptive admittance gains based on force-torque feedback, consistently produces lower interaction forces in all directions. The effectiveness of the CompACT compared to the baseline ACT was already presented in [17].
Figure 4: Force profiles of CompACT and ACT with GAC (fixed gains) during insertion tasks are presented. The CompACT with force-torque sensor inputs and output gains shows lower interaction force in all directions.
V-A2 Demonstration of Equivariance
Although the CompACT succeeds in insertion tasks in trained scenarios without excessive force exertion, it fails to generalize to spatially unseen configurations. This is expected, as its observation and action representations are defined in the global spatial frame, which neither guarantees nor encourages equivariance. The result of applying CompACT to the translationally unseen cases is shown in the Table I 33rd row - Flat Platform (OOD). Note that the flat platform is only utilized, meaning it is only randomized translationally. We tested for $10$ cases and did not conduct more tests because it resulted in a $0\%$ success rate.
In contrast, the proposed method (EquiContact) achieves perfect success rates on the translationally unseen flat platform, as can be seen in the $4$th row of Table I. As shown in Section IV, EquiContact has $SE(3)$ vision-to-force equivariance, achieving a near-perfect success rate, even on the tilted platform, which undergoes a full $SE(3)$ transformation. We attribute the single failure case to a large error from Diff-EDFs that exceeded the noise level applied during training.
TABLE II: Success Rates of the proposed EquiContact for a full pipeline of pick and place.
Test Scenario Success Rate Failure Cases
Flat Platform (OOD) 20 / 20 N/A
Tilted Platform ($30^{\circ}$, OOD) 18 / 20 1 Pick, 1 Place
The result of EquiContact for the full pick-and-place task is summarized in Table II. The EquiContact also demonstrates a near-perfect success rate in the full pick-and-place pipeline for peg-in-hole tasks. However, as the whole task is formulated in sequential stages, the error in the previous stage tends to propagate to the following stages, leading to slightly increased failure cases.
V-B Limitations and Discussions
TABLE III: Success Rates of the G-CompACT (insertion) trained with the base dataset and augmented dataset for the flat platform under the presence of visual distractor scenario and tilted platform with large angle cases. The evaluation is conducted with the ground-truth reference frames.
Test Scenario Base Dataset Aug. Dataset
W/ Visual Distractor 4 / 10 9 / 10
Tilted Platform ($45^{\circ}$) 5 / 10 9 / 10
V-B1 Limitation on Left-invariant Visual Features
Although EquiContact is theoretically vision-to-force equivariant, this guarantee holds only under Assumptions 1 and 2. In particular, Assumption 2, which requires left-invariant visual features, is challenging to ensure in practice, as our approach does not explicitly encode this property through loss functions or architectural inductive bias. Consequently, G-CompACT shows degraded performance in settings with visual distractors or severe tilting ($45^{\circ}$), as shown in Table III (“Base Dataset”). The degradation arises because larger tilt angles expose unseen background regions in the wrist camera view (see Fig. 3), resulting in out-of-distribution observations. This can be mitigated by augmenting the training data with visual distractors and demonstrations collected from a tilted angle. To demonstrate this, we added $20$ demos with distractors on a flat platform and 20 demos from a $30^{\circ}$ tilted setup. As shown in the “Aug. Dataset” column of Table III, the success rates have recovered to original levels.
Enforcing left-invariance on RGB images via architectural design remains an open challenge. Recent works have explored inducing 3D equivariance from 2D images [32, 33]. Another promising direction involves using pretrained foundation models, such as CLIP and DINO, which may exhibit implicit equivariance due to large-scale and diverse training. Language-guided features (e.g., using the prompt “peg”) can help extract relevant features across views. Alternatively, real-time segmentation models can be used to focus only on the object of interest. We will investigate enforcing left-invariance of images in future work.
V-B2 Hardware and Algorithmic Limitation
A key challenge in implementing equivariant policies for robotic manipulation is handling singularities [10]. When the robot is near singular, pose tracking accuracy degrades because controllers sacrifice tracking to avoid singularity, leading to poorly executed policy commands. This results in a distributional shift, which is a critical limitation for imitation learning. From our experimental demonstration perspective, a small workspace of the Diff-EDF (Fig. 2) was a drawback due to its limited receptive field. A combination of the singularity, joint limits, and small workspace resulted in limited demonstration of our approach, not covering a complete $SE(3)$ transformation.
V-C Validating Feasibility to Other Contact-rich Tasks
To further validate the EquiContact framework, we test it on two additional contact-rich tasks: screwing and surface wiping (see Fig. 5). In the screwing task, the robot aligns and screws a peg; in the wiping task, it erases a line from a whiteboard. For simplicity, we use a white background to eliminate visual distractions and assume that reference frames are known. The reference frame for screwing is the end-effector pose at full insertion; for wiping, it is the center of the board. As in the PiH setup, demonstrations are collected on a fixed platform and evaluated on out-of-distribution configurations, including tilted platforms. We collected 50 demonstrations per task. Results in Table IV show consistent success rates across all conditions, confirming EquiContact’s spatial generalizability.
Figure 5: (Left) Screwing task, (Right) Surface wiping (erasing) task. The same platform structure of the PiH is used.
TABLE IV: Success Rates of the G-CompACT for screwing and surface wiping tasks. The evaluation is conducted with the ground-truth reference frames. Both tasks are trained with $50$ demos.
Test Scenario Screwing Surface Wiping
Flat Platform (OOD) 10 / 10 10 / 10
Tilted Platform ($30^{\circ}$, OOD) 9 / 10 9 / 10
VI Conclusion
In this work, we introduced EquiContact, a vision-to-force equivariant policy for spatially generalizable contact-rich tasks. By integrating a global reference frame estimator (Diff-EDF) with a fully localized visuomotor servoing policy module (G-CompACT), we demonstrate how compliance, localized policy, and induced equivariance can be unified to enable the peg-in-hole (PiH) task, a representative contact-rich precision task, under spatial perturbations. We proved the $SE(3)$ equivariance of the policy under assumptions on point cloud and image observations, validated its effectiveness through real-world experiments on PiH benchmarks and its feasibility towards screwing and surface wiping tasks. Compared to benchmark methods, our approach generalizes to unseen platform positions and orientations while maintaining low contact force and near-perfect success rates. Through extensive benchmark studies, we highlighted the effectiveness of the three principles – compliance, localized policy, and induced equivariance – for achieving spatial generalizability in contact-rich manipulation. We conclude that these principles offer a simple yet powerful design guideline for developing spatially generalizable and interpretable robotic policies complementing recent trends in end-to-end visuomotor learning and enabling a structured divide-and-conquer approach.
Supplementary Material
Implementation Details
In this section, we provide the details of the hardware code implementation. Since every hardware system has a different implementation, our goal is to provide a comprehensive overview and a guide to our code distribution. We will use the exact names of the nodes to enable readers to understand our codebase easily. The hardware experimental implementation code is provided in https://github.com/Joohwan-Seo/EquiContact-Experiment, and the simulation implementation code in https://github.com/Joohwan-Seo/EquiContact-Simulation.
S-A Implementation Pipeline
The pipeline visualized in Fig. S1 is an interconnected system of camera, robot, and controller nodes with two policy services, all implemented in ROS2 Humble. The two policies are implemented as services because they are activated infrequently and latencies in activating the policies initially do not affect the behavior in practice. The execution of the full pipeline is divided into two primary stages: querying Diffusion-EDF for initial pick and insertion end-effector poses, and then executing the appropriate ACT sub-policies for these two steps after reaching a neighborhood of the predicted poses. In the first stage, a request to the Diffusion-EDF Service is made through the FullPipelineClient, which prompts the service to pull the individual point clouds published by the Orbbec Node. The service subsequently merges the point cloud in the spatial frame, preprocesses it, and passes it to the initialized Diff-EDF model to produce both pick and insertion poses. The Diff-EDF model, as covered in the following section, is formulated as a “double-pick” model that assumes that a fixed grip will be executed, and a cached empty gripper point cloud is passed along with the scene point cloud to get both predictions.
In the second stage, the FullPipelineClient requests the G-CompACT Service node to complete the task with these two poses as noisy references. The service initializes both pick and place policies and executes the tasks serially, raising flags to signal completion of each step. For raw camera images and proprioceptive information, the service subscribes to nodes for each camera (from Wrist Camera Nodes) and the Robot Node respectively; note that the controller and proprioceptive helper are encapsulated in one Robot Node that manages both publishing proprioceptive information and executing policy predictions. During the pick portion, the pick ACT model only publishes end-effector poses, and fixed high admittance gains are utilized because the demonstrations encourage the model to make limited contact with the peg. The pick model has a fixed time to complete the task before a grasp is attempted, and the robot moves to the next step. Similarly, the place ACT policy is executed given a fixed-time limit, but for the place model, the force-torque data is also passed as an input, and admittance gains are outputted along with the end-effector poses. A fixed-time limit for both phases is used because it is easily implemented and works more consistently than other, more complicated completion heuristics.
Figure S1: Whole pipeline implemented (via client named as FullPipelineClient) with ROS2 is presented.
S-B Diffusion-EDF
S-B1 Implementation Heuristics
Instead of following the original Diff-EDF pipeline, which utilizes pick-and-place models, we used two pick models. This decision mainly stems from the task setup, where the peg is upright, and the peg is grasped by the gripper in an aligned, upright pose. The core difference of the pick and place model is that the place model needs to get the grasp point cloud after each grasp to handle the right equivariance of the model. However, from task setup, we bypass this right equivariance issue, removing the necessity of the place model.
We also used the post-processing heuristics to filter the output pose of the Diff-EDF. The Diff-EDF outputs $20$ candidate target poses for picking and $10$ candidates for the place, which are ranked by the energy level. Although in theory, the lower energy poses should result in a better pose, we found out that this does not hold in practice. Instead, we figured out that the Diff-EDF have low variations on the position of the “tip”. To elaborate, suppose a picking policy returns the desired poses $g_{pick}^{i}=(p_{pick}^{i},R_{pick}^{i})$, which are the end-effector poses, not the tip, or Tool-Center-Point (TCP) poses. Then, the position of the tip estimated by Diff-EDF can be calculated via
$p_{tip,pick}^{i}=p_{pick}^{i}+R_{pick}^{i}p_{et,pick}$
where $p_{et,pick}\in\mathbb{R}^{3}$ denotes the coordinate of the gripper’s tip with respect to the end-effector body frame. Using this property and the task setup that the peg is always upright, which gives $R_{peg}$, we recalculate the Diff-EDF of the picking as
$p_{pick}=\bar{p}_{tip,pick}-R_{peg}^{T}p_{et,pick}$
where $\bar{p}_{tip,pick}$ is the mean of $p_{tip,pick}^{i}$ for $i=1,\cdots,N$. On the other hand, for placement, we did not make any assumptions about the orientation of the hole-platform structure. The heuristic we used for the place is as follows:
$p_{place}=\bar{p}_{tip,place}-\bar{R}_{place}^{T}p_{et,place},$
where $\bar{R}_{place}$ is the averaged rotation of $R_{place}^{i}$, and $\bar{p}_{tip,place}$ is the averaged position of the tip calculated similarly from the picking case.
S-C Tips on applying Geometric Admittance Control
Here, we present a few tips on the application of geometric admittance control. The GAC is implemented as a part of a ROS2 node, running with $200\mathrm{Hz}$.
S-C1 ROS2 Implementation
The Robot Node acts as the central controller, receiving and executing desired pose commands using a Geometric Admittance Control (GAC). Its main execution loop operates at $200\mathrm{Hz}$, ensuring smooth and compliant motion in response to external interactions. The node subscribes to the “action” topic, which contains the pose command and admittance gains command. Additionally, the controller subscribes to the “rebias” topic in order to preserve the accuracy of force readings. When a message on this topic is received, a callback raises a flag, triggering a re-biasing procedure for the FT sensor during the next control cycle. Within its main loop, the controller sends the intended pose and processed FT sensor data to the GAC function, which generates the compliant commands for the robot’s low-level motion controller.
S-C2 Handling Force-Torque Sensor Data
The first consideration is how to handle the Force-Torque (FT) sensor attached to the end-effector. It is always favorable to subtract the bias of the FT sensor after finishing each task, such as picking up the peg. We collect the FT sensor value for $2$ seconds and take a mean value to use as a bias. The other is to filter the FT sensor. Since the FT sensor is noisy, it is strongly recommended to implement a low-pass filter. However, having a low cut-off frequency can lead to serious phase lag, resulting in so-called “bouncing” behavior. We designed a simple 1st-order low-pass filter and implemented it using a state-space representation in discrete time.
S-C3 Implementation Tips
As mentioned in Section. IV, the detailed implementation method depends on the low-level servo controller type. Typical controllers provided by robot manufacturers are joint position control, joint velocity control, end-effector position control, and end-effector velocity control. If a torque controller is readily available, then directly implementing a geometric impedance controller (GIC) is recommended, but joint friction compensation needs to be handled appropriately. Given the desired dynamics (9), the desired end-effector pose command $\tilde{g}_{d}(k)$ provided to the end-effector controller is calculated in discrete time as
$V_{d}^{b}(k)\!$ $=\!V^{b}(k)\!+\!T_{s}\cdot M^{-1}(F_{e}(k)\!-\!f_{{}_{G}}(k)\!-\!K_{d}V^{b}(k)),$
$\tilde{g}_{d}(k)\!$ $=\!g(k)\cdot\exp{(\hat{V}_{d}^{b}(k)\cdot T_{s})},$ (16)
where $T_{s}$ is a sampling time ($5\mathrm{ms}$ in our case for GAC) and $\hat{(\cdot)}$ denotes a hat-map.
The end-effector velocity control can be straightforwardly applied from (S-C3), by directly using $V_{d}^{b}(k)$ as a desired end-effector velocity command. One caveat here is the $V^{b}$, which is body-frame end-effector velocity. The common velocity provided by the manufacturer is defined on the Cartesian (Spatial) frame. If end-effector velocity is only offered in the Cartesian frame $V\in\mathbb{R}^{6}$, then
$V^{b}=\begin{bmatrix}R^{T}&0_{3\times 3}\\ 0_{3\times 3}&R^{T}\end{bmatrix}V,$
where $R$ is a rotation matrix from the current configuration. However, sometimes, it is provided in the Cartesian frame for translational velocity and the body frame for rotational velocity. The users should verify the frames of the velocities by running the robot in one direction and logging its values. For the joint space-based approach, one can use
$\begin{split}\dot{q}_{d}(k)\!&=\!\dot{q}(k)\!+\!T_{s}\cdot J_{b}(k)^{-1}M^{-1}(F_{e}(k)\!-\!f_{{}_{G}}(k)\!-\!K_{d}V^{b}(k)),\\ q_{d}(k)\!&=\!q(k)+T_{s}\cdot\dot{q}_{d}(k),\end{split}$
where $J_{b}$ denotes a body-frame Jacobian matrix. $\dot{q}_{d}(k)$ is provided to joint-velocity controller and $q_{d}(k)$ is provided to the joint-position controller. The Jacobian matrix also needs to be identified concerning the end-effector velocity, i.e., if the end-effector velocity is given in the Cartesian frame, then the Jacobian matrix $J$ is also provided in the Cartesian frame. In this case, the body-frame Jacobian matrix can be obtained via
$J_{b}=\begin{bmatrix}R^{T}&0_{3\times 3}\\ 0_{3\times 3}&R^{T}\end{bmatrix}J$
When the Jacobian matrix is not provided, an alternative is to use the robotics toolbox from Python or Matlab. First, use the DH parameters provided by the manufacturer (or URDF), then build the Jacobian matrix function. The function can be cast into a lambda function in Python. For Python, the Jacobian matrix function can also be wrapped into executable .so file using Matlab code builder and executed in the loop for fast computation (well below $1\mathrm{ms}$).
S-D Teleoperation Details
The architecture of our teleoperation system is explained to make it easier to reproduce and adapt our work. The system is constructed modularly, with the human-operator interface decoupled from the robot controller. This allows greater adaptation to different robotic platforms and control algorithms. The teleoperation function is implemented by two ROS nodes: the $\mathtt{RobotNode}$ (named as $\mathtt{robot\_gac}$ in the code implementation), which processes proprioceptive information and executes a Geometric Admittance Controller to carry out operations, and the $\mathtt{Teleoperation}$ (named as $\mathtt{teleop\_ros}$ in the code implementation) node, which handles operator input.
The teleoperation node acts as the principal interface between the human operator and the robot controller. A 6-DOF SpaceMouse is used to record input, which is then converted into the necessary end-effector pose. The main teleop function of this node operates at a frequency of $200\mathrm{Hz}$ inside a timer loop. It adjusts the desired end-effector pose incrementally during each cycle by reading velocity commands from the SpaceMouse.
The SpaceMouse provides velocity commands $V_{sm}=[v_{sm}^{T},\omega_{sm}^{T}]^{T}\in\mathbb{R}^{6}$, where $v_{sm}\in\mathbb{R}^{3}$ is translational velocity and $\omega_{sm}\in\mathbb{R}^{3}$ is rotational velocity. The change in desired position, $\Delta p$, is computed by scaling $v_{sm}$ with a position offset:
$\Delta p=v_{sm}\cdot\mathtt{pos\_offset}$
The target position is then updated as:
$p_{d}(k+1)=p_{d}(k)+\Delta p$
The desired end-effector orientation, $R_{d}(k+1)$, is updated by composing the previous rotation with the incremental rotation matrix, $\Delta R$. The small, incremental angles, $\Delta\theta$, are computed by scaling $w_{sm}$ with rotation offset:
$\Delta\theta=w_{sm}\cdot\mathtt{rot\_offset}$
The $\Delta\theta$ is then converted to $\Delta R$ using Euler angle as $\Delta R=\mathtt{euler2rotmat(\Delta\theta)}$. The target orientation is then updated as:
$R_{d}(k+1)=R_{d}(k)\cdot\Delta R$
This calculated pose is then published as a command for the robot controller at a frequency of $200\mathrm{Hz}$. To create a dataset of the teleoperation commands, a separate timer logs the same message at $30\mathrm{Hz}$. This implements a teleoperation that operates on the robot base frame for translational movement and the end-effector frame for rotational movement.
This teleoperation node was created with multiple operation modes to improve usability. For end-effector alignment with tilted platforms, the operator can switch the SpaceMouse’s control frame between the end-effector frame and the robot’s base frame. Additionally, a dedicated button allows users to modify the sensitivity offsets for position and rotation, enabling users to alternate between a translation-focused and an orientation-focused mode. The operator can globally scale motion sensitivity by cycling through low, medium, and high speed levels. The system also includes specific features, such as an “insertion mode” that operates in the end-effector frame for the translational movement to help with high-precision alignment tasks, as well as requests to set the robot’s end effector at “home” or a randomized pose.
Furthermore, the operator can dynamically adjust the parameters of the Geometric Admittance Controller via dedicated keyboard inputs. There are four gain profiles: Free mode applies high stiffness gains $(1000)$ to all axes for precise motion in open space; Contact mode uses low z-axis gain $(300)$ with high gain elsewhere $(1500)$ to exert less force on surfaces; Insertion mode reverses this, with high z-axis gain $(1500)$ for precision while allowing compliance in other directions $(300)$; and Compliant mode sets all gains low $(300)$ to make the end-effector highly responsive to external forces. These gain profiles are published as action messages, together with poses and gripper states.
S-E G-CompACT Training Details
During the training of the G-CompACT, we added the noise to the known fixed pose of the hole and peg. As mentioned earlier, this is because the output of the Diff-EDFs $g_{{}_{EDF}}$ is biased. By adding the noise to the goal pose signal, we encourage the model to focus on the image of the wrist camera instead of the incorrect GCEV $e_{{}_{G}}$ signal. For our training, we added noise to every step of the demonstration dataset drawn from a uniform distribution. The noisy target poses $\tilde{g}_{ref}=(\tilde{p}_{ref},\tilde{R}_{ref})$are calculated as
$\begin{split}\tilde{p}_{ref}&=p_{ref}+n_{p},\\ \tilde{R}_{ref}&=R_{ref}\cdot\mathtt{euler2rotmat}(n_{r}),\end{split}$
where $n_{p}\sim\mathcal{U}(-0.02,0.02)\in\mathbb{R}^{3}$ and $n_{r}\sim\mathcal{U}(-8^{\circ},8^{\circ})\in\mathbb{R}^{3}$, which denotes $2\mathrm{cm}$ of translational and $8^{\circ}$ (degree) of rotational error. The impact of adding noise to the reference frame is further shown in Appendix. -C.
To encode the element on $SE(3)$ for learning, we used position & rotation vector representation for G-CompACT. However, in benchmark models of CompACT and ACTs, we utilized position & rot6d representation. This is because the rotation vector representation has a sign flip issue when the rotation angle is near $180^{\circ}$, but the rot6d representation does not have that problem. For G-CompACT, as the actions are represented with respect to the current end-effector frame, the transformation angles are relatively small and do not have a sign flip problem.